### 16 Apr 2026 v2

# Beatty solutions of almost Golomb equations

Abstract
The Golomb equation of order $r$ is the implicit functional equation
$a\Bigl(\sum_{j=0}^{r-1}a(n{-}j)\Bigr)=n$
for nondecreasing sequences of positive integers. Its earliest solution, the almost Golomb sequence of order $r$, is $r$-regular in the sense of Allouche and Shallit and has oscillating ratio $a(n)/n$. We prove that for every $r\geq 2$ that is not an even perfect square, the equation admits a second monotone solution given by an inhomogeneous Beatty sequence.
2020 Mathematics Subject Classification. Primary 11B83. Secondary 11B37, 39B12, 91A46.
Keywords. Beatty sequence, almost Golomb sequence, functional equation, Sturmian word, Wythoff game.
1 Introduction
1.1 The almost Golomb equation
Golomb’s sequence [5] is the unique nondecreasing sequence of positive integers in which $n$ appears exactly $a(n)$ times. It satisfies $a(\sum_{k=1}^{n}a(k))=n$ and grows like $a(n)\sim\varphi^{2-\varphi}n^{\varphi-1}$, where $\varphi=(1+\sqrt{5})/2$ is the golden ratio.
Replacing the cumulative sum by a sliding window of size $r$ leads to the Golomb equation of order $r$
$a\Bigl(\sum_{j=0}^{r-1}a(n{-}j)\Bigr)=n,\qquad n\geq 1,$ (1)
with $a(k)=0$ for $k<1$. The earliest nondecreasing solution, called the almost Golomb sequence of order $r$ in [2], is obtained inductively: $a(n)$ is the smallest integer $\geq a(n{-}1)$ consistent with (1). This finite-memory truncation changes the nature of the sequence. The smooth power law gives way to oscillatory linear growth, and the greedy solution is $r$-regular in the sense of Allouche and Shallit [1] for every $r\geq 2$ [2]. Writing $f(n)=\sum_{j=0}^{r-1}a(n{-}j)$, equation (1) reads $a\circ f=\mathrm{id}$.
Equation (1) is implicit. It constrains $a$ at position $f(n)$, which depends on $a$ itself and typically lies far beyond $n$. This self-referential addressing creates room for multiple solutions beyond the greedy one.
1.2 The greedy solution
The greedy (earliest) solution of (1) is studied in detail in the companion paper [2]. It is unique, $r$-regular, and has explicit $r$-adic denesting formulas. For $r=2$ (A394217 in [8]), it begins
[…]
and satisfies the dyadic recurrences a(2n)=a(n)+a(n+1)1a(2n)=a(n)+a(n{+}1)-1 and a(2n+1)=a(n)+a(n+1)a(2n{+}1)=a(n)+a(n{+}1) [2]. Its ratio $a(n)/n$ oscillates between $2/3$ and $3/4$ and does not converge [2].
1.3 A second solution
In this paper we show that, once the greedy constraint is removed, the same implicit equation admits a second monotone solution of a different nature. The case $r=2$ already illustrates the phenomenon.
Theorem 1.
The inhomogeneous Beatty sequence
$a(n)=\Bigl\lfloor\frac{n+1}{\sqrt{2}}\Bigr\rfloor,\qquad n\geq 1,$ (2)
with $a(k)=0$ for $k<1$, satisfies the Golomb equation of order $2$:
$a\bigl(a(n)+a(n{-}1)\bigr)=n\qquad\text{for all }n\geq 1.$ (3)
This Beatty solution begins
[…]
The two solutions agree on $[1,11]$ and diverge at $n=12$, where the greedy repeats the value $8$ (run of length $2$), while the Beatty moves to $9$ (run of length $1$). This is possible because the constraint at $n=12$ applies at position $f(12)=a(12)+a(11)$, which is $16$ for the greedy and $17$ for the Beatty. Different values of $a$ shift the constrained positions, and both are self-consistent.
In the rest of the paper, $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$ denote the greedy and the canonical Beatty solutions for the case $r=2$. Otherwise $a$ is generic and is made precise by the context. The table below compares the two solutions and their constrained positions $S_{n}=a(n)+a(n{-}1)$ near the first divergence point.
$n$ $a_{\mathrm{G}}(n)$ $a_{\mathrm{B}}(n)$ $S^{\mathrm{G}}_{n}$ $S^{\mathrm{B}}_{n}$ $a_{\mathrm{G}}(S^{\mathrm{G}}_{n})$ $a_{\mathrm{B}}(S^{\mathrm{B}}_{n})$
10 7 7 14 14 10 10
11 8 8 15 15 11 11
12 8 9 16 17 12 12
13 9 9 17 18 13 13
14 10 10 19 19 14 14
Greedy Beatty
Rule local, minimal global, irrational slope
Structure 2-regular, dyadic Sturmian, Ostrowski
$a(n)/n$ oscillates ($2/3\leftrightarrow 3/4$) converges to $1/\!\sqrt{2}$
Multiplicity 22-automatic Sturmian
Figure 1 shows both solutions over the range $n\leq 512$. The greedy $a_{\mathrm{G}}$ crosses the Beatty $a_{\mathrm{B}}$ repeatedly: the residual $a_{\mathrm{B}}(n)-n/\!\sqrt{2}$ is bounded (it is the Sturmian remainder $\{(n{+}1)/\!\sqrt{2}\}$), whereas $a_{\mathrm{G}}(n)-n/\!\sqrt{2}$ grows without bound because $a_{\mathrm{G}}(n)/n$ oscillates in $[2/3,3/4]$ rather than converging to $1/\!\sqrt{2}$. Figure 2 below illustrates the difference between $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$.
Figure 1: The greedy solution $a_{\mathrm{G}}(n)$ and the canonical Beatty $a_{\mathrm{B}}(n)$ for $1\leq n\leq 512$.
Figure 2: The normalised difference $(a_{\mathrm{G}}(n)-a_{\mathrm{B}}(n))/n$ for $1\leq n\leq 512$.
Theorem 1 is the case $r=2$ of the following general result.
Theorem 2.
Let $r\geq 2$ and define the order-$r$ Beatty sequence
$a(n)=\Bigl\lfloor\frac{n}{\sqrt{r}}+\frac{\sqrt{r}}{2}\Bigr\rfloor.$ (4)
Then $a\bigl(\sum_{j=0}^{r-1}a(n{-}j)\bigr)=n$ for all $n\geq r$ if and only if $r$ is not an even perfect square.
When $r=(2m)^{2}$, the left-hand side equals $n+1$ instead of $n$ for all $n\geq r$ (Theorem 12).
The proof combines three ingredients. The continuous functional equation identifies the unique affine candidate (Section 2). Floor discretisation then reduces the identity to a finite counting inequality $N(r)<r\beta+s-1$, where $N(r)$ counts certain fractional parts of multiples of $1/\!\sqrt{r}$ (Proposition 9). A parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$ establishes this inequality for all non-square $r$ (Section 3).
1.4 The triple-nested family
Applying $a$ to both sides of (3) gives $a(a(f(n)))=a(n)$, the triple-nested equation
$a\bigl(a(a(n)+a(n{-}1))\bigr)=a(n).$ (5)
The identity (3) selects a unique value of $d$ (namely $d=\sqrt{2}/2$), but the weaker identity (5) admits a continuous family.
Theorem 3.
Let $c=1/\!\sqrt{2}$ and I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}-1)]. For every $d\in I$, the sequence defined by $a(n)=\lfloor cn+d\rfloor$ satisfies (5) for all $n\geq 1$. The interval $I$ is sharp. For $d\notin I$ the identity fails at infinitely many $n$.
The endpoints of $I$ are the two Beatty parameters associated with the $\sqrt{2}$-Wythoff game (Fraenkel [4]).
1.5 Outline
Sections 23 prove the existence results. Section 2 establishes the exact continuous solution. Section 3 reduces the discrete identity to a counting inequality via sawtooth analysis, then closes it by a parity argument for all $r$ that are not even perfect squares. Section 4 proves the Beatty interval for the triple-nested equation (Theorem 3). The remaining sections discuss structural phenomena specific to $r=2$: Section 5 discusses the Sturmian structure and bifurcation behaviour in the case $r=2$, and Section 6 studies a conjectural defect morphism at the boundary of the Beatty interval.
2 The continuous equation
In the discrete equation (1), $f(n)=\sum a(n{-}j)$ denotes the sliding-window sum and the identity reads $a\circ f=\mathrm{id}$. The continuous analogue replaces $a$ by a smooth function $\varphi$:
$\varphi\Bigl(\sum_{j=0}^{r-1}\varphi(x{-}j)\Bigr)=x.$ (6)
The affine function
$\varphi(x)=\frac{x}{\sqrt{r}}+\frac{\sqrt{r}-1}{2}$ (7)
satisfies (6), as one checks by substitution.
Theorem 4.
The function (7) is the unique affine solution of (6). Moreover, every continuous nondecreasing solution is a homeomorphism of $\mathbb{R}$ satisfying the inverse equation
$\varphi^{-1}(x)=\sum_{j=0}^{r-1}\varphi(x{-}j).$ (8)
Proof.
Affine solutions. If $\varphi(x)=cx+d$ then
$\sum_{j=0}^{r-1}\varphi(x{-}j)=r(cx+d)-\frac{cr(r{-}1)}{2}$
and then
$\varphi\Bigl(\sum\varphi(x{-}j)\Bigr)=rc^{2}x+\Bigl(rcd-\frac{rc^{2}(r{-}1)}{2}+d\Bigr).$
This equals $x$ if and only if $rc^{2}=1$ and $d(\sqrt{r}+1)=(r{-}1)/2$. The unique positive solution is $c=1/\!\sqrt{r}$, $d=(\sqrt{r}-1)/2$.
Rigidity. Let $\varphi$ be any continuous nondecreasing solution and set $g(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$. Then $g$ is continuous and nondecreasing, and $\varphi(g(x))=x$ for all $x$, so $\varphi$ is surjective and $g$ is injective. Since $g$ is continuous, nondecreasing, and injective, it is strictly increasing. If $\varphi$ had a flat interval $[u,v]$ with $\varphi\equiv\lambda$ there, then $g(\lambda)\in[u,v]$ but $g(x)<u$ for $x<\lambda$ and $g(x)>v$ for $x>\lambda$, contradicting continuity of $g$ at $\lambda$. So $\varphi$ is strictly increasing, hence a homeomorphism, and $g=\varphi^{-1}$. ∎
Remark 5.
The discrete Beatty solution $a(n)=\lfloor n/\!\sqrt{r}+\sqrt{r}/2\rfloor$ is the floor-discretisation of $\varphi$: the shift $d_{\mathrm{Beatty}}=\sqrt{r}/2$ equals $d_{\mathrm{cont}}+1/2=(\sqrt{r}-1)/2+1/2$, the universal +1/2+1/2 correcting for the average of the fractional part.
Whether the inverse equation (8) already forces $\varphi$ to be affine remains open (see Section 8).
3 The universal identity
Throughout this section we fix $r\geq 2$ and write $a$ for the order-$r$ Beatty sequence
$a(n)=\Bigl\lfloor\frac{n}{\alpha}+\frac{\alpha}{2}\Bigr\rfloor,$ (9)
where $\alpha=\sqrt{r}$ and $c=1/\alpha$. Set
$S(n)=\sum_{j=0}^{r-1}a(n{-}j).$
3.1 The discretisation residual
By Theorem 4, the continuous identity $\varphi(\sum\varphi(x{-}j))=x$ is exact. The only obstruction to the discrete identity $a(S(n))=n$ is the rounding error introduced by the floor. Define
$\Phi(n):=\sum_{j=0}^{r-1}\Bigl\{\frac{n{-}j}{\alpha}+\frac{\alpha}{2}\Bigr\}.$
Proposition 6.
For all $n\geq r$,
$a(S(n))=n\quad\iff\quad\frac{r-\alpha}{2}<\Phi(n)\leq\frac{r+\alpha}{2}.$
Proof.
Write $x_{j}=(n{-}j)/\alpha+\alpha/2$, so that $a(n{-}j)=x_{j}-\{x_{j}\}$. Summing over $j=0,\ldots,r{-}1$:
$S(n)=\sum x_{j}-\Phi(n).$
Since $\alpha^{2}=r$ we get $\sum x_{j}=(n+1/2)\alpha$, hence $S(n)=(n+1/2)\alpha-\Phi(n)$. Applying $a$:
$a(S(n))=\Bigl\lfloor n+\frac{1+\alpha}{2}-\frac{\Phi(n)}{\alpha}\Bigr\rfloor.$
This equals $n$ iff $0\leq(1{+}\alpha)/2-\Phi(n)/\alpha<1$, which rearranges to $(r{-}\alpha)/2<\Phi(n)\leq(r{+}\alpha)/2$. ∎
3.2 Counting reformulation
Write $s=\lfloor\alpha\rfloor$, $\beta=\alpha-s\in(0,1)$, and let
$N(r)=\#\bigl\{1\leq j\leq r{-}1:\{j/\alpha\}<\beta\bigr\}.$ (10)
Lemma 7.
The pairing $j\leftrightarrow r{-}j$ gives
$2\sum_{j=1}^{r-1}\lfloor jc\rfloor=(r{-}1)(s{-}1)+N(r),$ (11)
and $N(r)<r\beta+s-1$ is equivalent to
$\sum_{j=1}^{r-1}\{jc\}>\frac{r-\alpha}{2}.$ (12)
Proof.
Since $jc+(r{-}j)c=\alpha\notin\mathbb{Z}$, $\lfloor jc\rfloor+\lfloor(r{-}j)c\rfloor=s$ when $\{jc\}<\beta$ and =s1=s{-}1 otherwise, giving (11). Writing $\sum\{jc\}=(r{-}1)\alpha/2-[(r{-}1)(s{-}1)+N(r)]/2$, the equivalence with (12) is algebraic. ∎
The local maximum of the sawtooth just before the jump at $\theta=\{j_{0}c\}$ is
$M(j_{0})=r-j_{0}+\sum_{k=1}^{j_{0}}\{kc\}-\sum_{m=1}^{r-1-j_{0}}\{mc\},$ (13)
and the local minimum just after the same jump is $M(j_{0})-1$ (using $\{-mc\}=1-\{mc\}$ for $m\geq 1$). Figure 3 shows the sawtooth function over a full period for $r=2,3,5$.
Figure 3: The sawtooth $\Phi_{r}(\theta)=\sum_{j=0}^{r-1}\{\theta-jc\}$ for $r=2,3,5$ with $c=1/\!\sqrt{r}$, over one period $\theta\in[0,1)$.
The supremum and infimum of the sawtooth $\Phi$ enjoy a palindromic symmetry.
Lemma 8.
$\sup\Phi+\inf\Phi=r$.
Proof.
Set $\Delta_{k}=\{kc\}+\{(r{-}k)c\}-1$ for $k=1,\ldots,r{-}1$ and $P(j)=\sum_{k=1}^{j}\Delta_{k}$, with $P(0)=0$. By (13), $M(j_{0})=r-\Sigma+P(j_{0})$ where $\Sigma=\sum_{j=1}^{r-1}\{jc\}$. Since $\{kc\}+\{(r{-}k)c\}$ is symmetric in $k$ and rkr{-}k, $\Delta_{k}=\Delta_{r-k}$, so the step sequence is palindromic and P(j)+P(r1j)=P(r1)P(j)+P(r{-}1{-}j)=P(r{-}1) for all $j$. Therefore $\max P+\min P=P(r{-}1)$. Then
$\sup\Phi+\inf\Phi$ $=(r{-}\Sigma{+}\max P)+(r{-}1{-}\Sigma{+}\min P)$
$=2r{-}1{-}2\Sigma+P(r{-}1).$
Substituting $P(r{-}1)=(r{-}1)\beta-N(r)$ and $2\Sigma=(r{-}1)(\beta{+}1)-N(r)$ (from (11)) gives $2r{-}1{-}2\Sigma+P(r{-}1)=r$. ∎
Proposition 9.
For non-square $r$, $N(r)<r\beta+s-1$ implies $a(S(n))=n$ for all $n\geq r$.
Proof.
By Proposition 6 and the density of $\{cn+\alpha/2\}$ in $[0,1)$, the identity holds iff $(r{-}\alpha)/2<\Phi(\theta)\leq(r{+}\alpha)/2$ for all $\theta$. By Lemma 8, $\sup\Phi+\inf\Phi=r$, so the two bounds are equivalent: it suffices to show $\sup\Phi<(r{+}\alpha)/2$ (equality excluded by irrationality of $c$).
By (13), $M(j_{0})=r-\Sigma+P(j_{0})$ where $\Sigma=\sum\{jc\}$ and $P(j)=\sum_{k=1}^{j}\Delta_{k}$ as in Lemma 8. So $\sup\Phi=r-\Sigma+\max P$, and the required bound becomes $\max P<\Sigma-(r{-}\alpha)/2$.
Since each step $\Delta_{k}$ lies in $\{\beta{-}1,\,\beta\}$, the partial sum $P(j)=j\beta-C(j)$ where $C(j)=\#\{1\leq k\leq j:\{kc\}<\beta\}$. The sequence $\{kc\}$ is a Sturmian rotation, and the bounded remainder property of irrational rotations gives $|C(j)-j\beta|<1$ for all $j$ (this is the balance property of Sturmian words, see [7, Chapter 2]). Hence $|P(j)|<1$ for every $j$. By Lemma 8, $\max P+\min P=P(r{-}1)$, so $\max P-\min P<2$.
Now $\sup\Phi<(r{+}\alpha)/2$ is equivalent to $\max P-\min P<\alpha{-}1$ (substituting $\Sigma-(r{-}\alpha)/2=[\alpha{-}1{-}P(r{-}1)]/2$ and using $\max P-\min P=2\max P-P(r{-}1)$). For non-square $r\geq 10$ one has $\alpha{-}1>2$, so the bound $\max P-\min P<2$ suffices. For $3\leq r\leq 8$ non-square, the bound is checked by direct evaluation of the relevant quantities, as recorded in Table 1. For $r=2$, $\sup\Phi=(r{+}\alpha)/2$ exactly, but $c\notin\mathbb{Q}$ ensures the supremum is not attained. ∎
The quantities entering the counting inequality $N(r)<r\beta+s-1$ for the small cases $r\in\{2,3,5,6,7,8\}$ are collected below, together with the resulting extrema $\sup\Phi$ and $\inf\Phi$. The margin column $r\beta+s-1-N(r)$ is positive in every case.
$r$ $s$ $\beta$ $N(r)$ $r\beta+s-1$ margin $\sup\Phi$ $\inf\Phi$
22 11 $0.4142$ 0 $0.8284$ $0.8284$ $1.7071$ $0.2929$
33 11 $0.7321$ 22 $2.1962$ $0.1962$ $2.2679$ $0.7321$
55 22 $0.2361$ 0 $2.1803$ $2.1803$ $3.4721$ $1.5279$
66 22 $0.4495$ 33 $3.6969$ $0.6969$ $3.8763$ $2.1237$
77 22 $0.6458$ 44 $5.5203$ $1.5203$ $4.3542$ $2.6458$
88 22 $0.8284$ 77 $7.6274$ $0.6274$ $5.1005$ $2.8995$
Table 1: Direct verification of the counting inequality $N(r)<r\beta+s-1$ for small non-square $r$. The margin column is $r\beta+s-1-N(r)$, positive in every case. The extrema satisfy $\sup\Phi+\inf\Phi=r$ (Lemma 8) and both lie in the required interval $((r-\alpha)/2,(r+\alpha)/2]$. The tightest case is $r=3$ with margin $\approx 0.196$.
3.3 The parity argument
Write $r=s^{2}+t$ with $1\leq t\leq 2s$, so that $\alpha=s+\beta$ with $\beta\in(0,1)$ and $\beta(2s+\beta)=t$. The key observation is that this relation allows us to trade the irrational parameter $\beta$ for the integer $t$, whose parity determines the floor $u=\lfloor s\beta\rfloor$ and thereby the structure of the counting bound.
Lemma 10.
$\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$.
Proof.
Since $\beta=t/(s+\alpha)$ and $s+\alpha\in(2s,2s{+}1)$, we get $t/(2s{+}1)<\beta<t/(2s)$. Multiplying by $s$: $st/(2s{+}1)<s\beta<t/2$. Since st/(2s+1)(t1)/2=(2st+1)/$(2$s+1))>0st/(2s{+}1)-(t{-}1)/2=(2s{-}t{+}1)/$(2$s{+}1))>0 (as $t\leq 2s$), we obtain $(t{-}1)/2<s\beta<t/2$, whence $\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$. ∎
Lemma 11.
$N(r)=(s{+}1)(t{-}1)-2\sum_{k=1}^{s}\lfloor k\beta\rfloor$.
Proof.
For $0\leq m\leq s{-}1$, the number of integers $j\in[1,r{-}1]$ with $\lfloor j/\alpha\rfloor=m$ and $\{j/\alpha\}<\beta$ is $C_{m}=t+\lfloor(m{-}s)\beta\rfloor-\lfloor m\beta\rfloor$ (since $m\alpha+\alpha\beta=m\alpha+t-s\beta$ and the endpoints are irrational). For $m=s$: $C_{s}=t{-}1{-}\lfloor s\beta\rfloor$. Using $\lfloor-m\beta\rfloor=-\lfloor m\beta\rfloor-1$ for $m\geq 1$ and summing gives the result. ∎
Since $\lfloor k\beta\rfloor$ counts integers $m\geq 1$ with $m\leq k\beta$, swapping the order of summation gives
$\sum_{k=1}^{s}\lfloor k\beta\rfloor=\sum_{m=1}^{u}\bigl(s{+}1-\lceil m/\beta\rceil\bigr),$ (14)
where $u=\lfloor s\beta\rfloor$.
We now prove Theorem 2.
Proof of Theorem 2.
By Proposition 9, it suffices to show $N(r)<r\beta+s-1$. By Lemmas 10 and $1$, we split by the parity of $t$.
Case t=2u+1t=2u{+}1 odd. Then $\lfloor s\beta\rfloor=u$, and substituting (14) into Lemma 11 gives $N(r)=2\sum_{m=1}^{u}\lceil m/\beta\rceil$. Since $\beta$ is irrational, each $\lceil m/\beta\rceil<m/\beta+1$, so
$N(r)<\frac{u(u{+}1)}{\beta}+2u.$
The margin is
$r\beta+s-1-\frac{u(u{+}1)}{\beta}-2u=\frac{(1{-}\beta)^{2}(\beta^{2}{+}2\beta{+}4u{+}3)}{4\beta}>0$
(the equality is checked by substituting $r=s^{2}+t$ and $\beta(2s{+}\beta)=t=2u{+}1$, then clearing denominators).
Case $t=2u$ even. Then $\lfloor s\beta\rfloor=u{-}1$, and the same substitution gives $N(r)=s{+}1+2\sum_{m=1}^{u-1}\lceil m/\beta\rceil$. Using $\lceil m/\beta\rceil<m/\beta+1$ again:
$N(r)<s+1+\frac{u(u{-}1)}{\beta}+2u-2.$
The margin is
$r\beta+s-1-s-1-\frac{u(u{-}1)}{\beta}-2u+2=\frac{u(1{-}\beta)^{2}}{\beta}+\frac{\beta^{3}}{4}>0$
(checked by substituting $\beta(2s{+}\beta)=t=2u$ and clearing denominators).
In both cases $N(r)<r\beta+s{-}1$ and the theorem follows from Proposition 9. ∎
3.4 Perfect squares
When $r=s^{2}$, the parameter $\beta$ vanishes and the counting argument of the previous subsection degenerates. The window sum can instead be evaluated exactly by grouping terms modulo $s$, which leads to a clean application of Hermite’s identity.
Theorem 12.
Let $r=s^{2}$ and $a(n)=\lfloor n/s+s/2\rfloor$. For all $n\geq s^{2}$:
  1. 1.
    if $s$ is odd, S(n)=sns(s1)/2S(n)=sn-s(s{-}1)/2 and $a(S(n))=n$;
  2. 2.
    if $s$ is even, S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2 and $a(S(n))=n+1$.
Proof.
Write $q=\lfloor s/2\rfloor$.
Case $s$ odd (s=2q+1s=2q{+}1). Then $d=s/2=q+1/2\notin\mathbb{Z}$, and $a(n)=q+\lfloor(n{+}q)/s\rfloor$. For $n\geq s^{2}$ the full window is active, so
$S(n)=s^{2}q+\sum_{j=0}^{s^{2}-1}\Bigl\lfloor\frac{n-j+q}{s}\Bigr\rfloor.$
Write $j=ks+u$ with $0\leq k,u\leq s{-}1$. Then $\lfloor(n{-}ks{-}u{+}q)/s\rfloor=\lfloor(n{-}u{+}q)/s\rfloor-k$, giving
$S(n)=s^{2}q+s\sum_{u=0}^{s-1}\Bigl\lfloor\frac{n{-}u{+}q}{s}\Bigr\rfloor-s\cdot\frac{s(s{-}1)}{2}.$
As $u$ ranges over $\{0,\ldots,s{-}1\}$, the values n+qun{+}q{-}u hit each residue class modulo $s$ exactly once, so by Hermite’s identity ($\sum_{v=0}^{n-1}\lfloor x+v/n\rfloor=\lfloor nx\rfloor$ for all $x\in\mathbb{R}$, see [6, §3.5])
$\sum_{u=0}^{s-1}\Bigl\lfloor\frac{n{+}q{-}u}{s}\Bigr\rfloor=n+q-s+1.$
Substituting and using q=(s1)/2q=(s{-}1)/2 yields S(n)=sns(s1)/2S(n)=sn-s(s{-}1)/2. Finally, $a(S(n))=q+\lfloor n{-}q+q/s\rfloor=q+(n{-}q)=n$, since $0\leq q/s=(s{-}1)/(2s)<1$.
Case $s$ even ($s=2m$). Then $d=m\in\mathbb{Z}$ and $a(n)=m+\lfloor n/s\rfloor$. Applying Hermite’s identity in the same way gives S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2. Then $a(S(n))=m+\lfloor n{-}(s{-}2)/2\rfloor=n+1$, since $(s{-}2)/2=m{-}1\in\mathbb{Z}$. ∎
Corollary 13.
The canonical Beatty identity $a(S(n))=n$ holds for all odd perfect squares $r=s^{2}$ ($s=3,5,7,\ldots$) and fails systematically for all even perfect squares $r=(2m)^{2}$.
3.5 Coexistence of solutions
For each $r$ (excluding $r=(2m)^{2}$), the Golomb equation (1) of order $r$ admits at least two different monotone solutions: the greedy [2], which is $r$-regular with oscillating ratio $a(n)/n$, and the Beatty, with $a(n)/n\to 1/\!\sqrt{r}$. Both are fixed points of the implicit equation $a\circ f=\mathrm{id}$, sustained by different self-consistent address structures.
Remark 14.
The greedy branch is governed by finite-state local structure and is therefore amenable to $r$-regular methods [2, 1]. The Beatty branch is governed by irrational rotation and the continuous functional equation (6).
4 The triple-nested equation
We now compose the Golomb equation (1) of order $r$ with $a$, obtaining the weaker identity $a(a(f(n)))=a(n)$. This reveals a continuous family of Beatty solutions parametrised by a shift $d$ ranging over an explicit interval, with connections to Wythoff games and combinatorics on words.
4.1 The continuous perspective
For the strong identity $\varphi(g(x))=x$ and the triple-nested $\varphi(\varphi(g(x)))=\varphi(x)$ (where $g(x)=\sum\varphi(x{-}j)$), the continuous equation gives the same unique affine solution (7) in both cases: if $\varphi(g(x))=x$, composing with $\varphi$ immediately yields $\varphi(\varphi(g(x)))=\varphi(x)$, and conversely, injectivity of $\varphi$ recovers the strong identity from the triple-nested one.
In the discrete setting, the floor breaks this equivalence. The strong identity pins down a unique shift $d=\sqrt{r}/2$ (the left endpoint of the Beatty interval), while the weaker triple-nested identity tolerates a range of shifts. The floor in the outer composition $a(\cdot)$ absorbs small errors, opening the interval $I$ to the right.
4.2 Expansion of the triple composition
Throughout this section, $c=1/\!\sqrt{2}$, $d\in\mathbb{R}$, and $a(n)=\lfloor cn+d\rfloor$. Write $e(n)=-\{cn+d\}\in(-1,0]$, so that $a(n)=cn+d+e(n)$. Define f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1). Then
f(n)=c(2n1)+2d+e(n)+e(n1).f(n)=c(2n{-}1)+2d+e(n)+e(n{-}1). (15)
Applying $a$ to $f(n)$ and using $2c^{2}=1$:
$a(f(n))=\frac{2n{-}1}{2}+d(2c{+}1)+c\bigl(e(n)+e(n{-}1)\bigr)+e(f(n)).$ (16)
A second application of $a$ yields $a(a(f(n)))=cn+R(n)$, where
$R(n)=R_{0}+c^{2}(e(n)+e(n{-}1))+c\,e(f(n))+e(a(f(n))),$ (17)
with constant part
$R_{0}=d(c+2)-\frac{c}{2}.$ (18)
The equation (5) holds iff $\lfloor cn+R(n)\rfloor=\lfloor cn+d\rfloor$ for all $n\geq 1$.
4.3 Necessity
Any solution must satisfy $a(1)=\lfloor c+d\rfloor=1$, requiring $1-c\leq d<2-c$. The failure of (5) at both extremes of $I$ is detected along the continued-fraction convergents of $c$. We record the needed Pell data.
Lemma 15.
Let $p_{k}/q_{k}$ be the $k$-th convergent of $c=[0;1,2,2,2,\ldots]$ and set $\varepsilon_{k}=(-1)^{k}(cq_{k}-p_{k})>0$. The first values are
$k$ 11 22 33 44 55 66 77 88
$p_{k}$ 11 22 55 $12$ $29$ $70$ $169$ $408$
$q_{k}$ 11 33 77 $17$ $41$ $99$ $239$ $577$
$\varepsilon_{k}$ .293.293 .121.121 .050.050 .021.021 .009.009 .004.004 .001.001 .0006.0006
and the following identities hold for every $k\geq 1$:
  1. 1.
    $cq_{k}=p_{k}+(-1)^{k}\varepsilon_{k}$ with $\varepsilon_{k}\to 0^{+}$;
  2. 2.
    $2cp_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k}$.
Proof.
The numerical table is standard. Identity (i) is the definition of $\varepsilon_{k}$. Identity (ii) follows from (i): $2cp_{k}=2c(cq_{k}-(-1)^{k}\varepsilon_{k})=2c^{2}q_{k}-2c(-1)^{k}\varepsilon_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k}$, using $2c^{2}=1$. ∎
Lemma 16.
If $d<\sqrt{2}/2$, then (5) fails at infinitely many $n$.
Proof.
Write $d=c-\delta$ with $\delta>0$. Fix $k\geq 2$ even large enough that $\varepsilon_{k}<\delta$, so that (i) gives $cq_{k}=p_{k}+\varepsilon_{k}$ with $0<\varepsilon_{k}<\delta<c<1$. Evaluate at $n=q_{k}$:
$a(q_{k})$ $=\lfloor cq_{k}+d\rfloor=\lfloor p_{k}+\varepsilon_{k}+c-\delta\rfloor=p_{k},$
$a(q_{k}{-}1)$ $=\lfloor c(q_{k}{-}1)+d\rfloor=\lfloor p_{k}+\varepsilon_{k}-\delta\rfloor=p_{k}-1,$
where the first floor uses $\varepsilon_{k}+c-\delta\in(0,c)\subset(0,1)$ and the second uses $1<\varepsilon_{k}-\delta<0$. Hence $f(q_{k})=a(q_{k})+a(q_{k}{-}1)=2p_{k}-1$, and by identity (ii):
$a(f(q_{k}))=\lfloor c(2p_{k}-1)+d\rfloor=\lfloor q_{k}-2c\varepsilon_{k}-\delta\rfloor=q_{k}-1,$
since $0<2c\varepsilon_{k}+\delta<1$ for $k$ large. Applying $a$ once more:
$a(a(f(q_{k})))=a(q_{k}-1)=p_{k}-1\neq p_{k}=a(q_{k}),$
which violates (5) at $n=q_{k}$. Since infinitely many even $k$ are admissible, (5) fails at infinitely many $n$. ∎
Lemma 17.
If $d>2(\sqrt{2}-1)$, then (5) fails at infinitely many $n$.
Proof.
Write $u=4c-2=2(\sqrt{2}-1)$ and $d=u+\delta$ with $\delta>0$. The bound d<2cd<2-c forces $\delta<1-c$. Fix $k\geq 3$ odd large enough that $\varepsilon_{k}<\delta$ and $2c\varepsilon_{k}+\delta<1-c$, and take $n=q_{k}-3$. Identity (i) with $k$ odd gives $cq_{k}=p_{k}-\varepsilon_{k}$, hence
$a(q_{k}{-}3)$ $=\lfloor c(q_{k}{-}3)+d\rfloor=\lfloor p_{k}-\varepsilon_{k}-3c+u+\delta\rfloor$
$=\lfloor p_{k}-(2-c)+(\delta-\varepsilon_{k})\rfloor=p_{k}-2,$
since $2-c\approx 1.293$ and $0<\delta-\varepsilon_{k}<1-c$, so $(2-c)+(\delta-\varepsilon_{k})$ lies in $(c-2,c-1)\subset(-2,-1)$, giving floor value $p_{k}-2$. Similarly
$a(q_{k}{-}4)=\lfloor p_{k}-(2-2c)+(\delta-\varepsilon_{k})\rfloor=p_{k}-2,$
since $(2-2c)+(\delta-\varepsilon_{k})$ lies in $(2c-2,2c-1)\subset(-1,0)$. Hence $q_{k}-3$ is a repeat and $f(q_{k}{-}3)=2p_{k}-4$. By identity (ii),
$a(2p_{k}{-}4)$ $=\lfloor c(2p_{k}{-}4)+d\rfloor=\lfloor q_{k}+2c\varepsilon_{k}-4c+u+\delta\rfloor$
$=\lfloor q_{k}-(2-2c\varepsilon_{k}-\delta)\rfloor=q_{k}-2,$
since $2c\varepsilon_{k}+\delta<1-c<1$, so $(2-2c\varepsilon_{k}-\delta)\in(-2,-1)$. Finally
$a(q_{k}{-}2)=\lfloor p_{k}-\varepsilon_{k}-2c+u+\delta\rfloor=\lfloor p_{k}+(2c-2)+(\delta-\varepsilon_{k})\rfloor=p_{k}-1,$
since $(2c-2)+(\delta-\varepsilon_{k})\in(2c-2,1-2c)\subset(-1,0)$. Therefore
$a(a(f(q_{k}{-}3)))=a(q_{k}{-}2)=p_{k}-1\neq p_{k}-2=a(q_{k}{-}3),$
which violates (5) at $n=q_{k}-3$. Infinitely many odd $k$ are admissible. ∎
4.4 Sufficiency
Proof of Theorem 3.
Write $\theta=\{cn+d\}$, so $a(n)=cn+d-\theta$. The index $n$ is a repeat if $\theta<1{-}c$ (giving a(n+1)=a(n)a(n{+}1)=a(n)) and a non-repeat if $\theta\geq 1{-}c$. Set D=d(2c+1)D=d(2c{+}1).
Non-repeat, $\theta\in[c,1)$. Then $f(n)=2a(n)$ and $a(f(n))=\lfloor n+\varphi\rfloor$ with $\varphi=D-2c\theta\in(D-2c,\;D-1]$. Since $d\geq c$ gives $\varphi>0$ and $d\leq 4c{-}2$ gives $\varphi\leq 1$ with equality only when d=4c2d=4c{-}2 and $\theta=c$ (requiring $c(n{+}3)\in\mathbb{Z}$, impossible), $\varphi\in(0,1)$ and $a(f(n))=n$.
Non-repeat, $\theta\in[1{-}c,c)$. Then f(n)=2a(n)1f(n)=2a(n)-1 and $\varphi=D-c(2\theta{+}1)\in(0,1)$ by the same bounds on $D$.
Repeat ($\theta\in[0,1{-}c)$). Then $\varphi\in(0,2)$, so $a(f(n))\in\{n,n{+}1\}$. Since $n$ is a repeat, a(n+1)=a(n)a(n{+}1)=a(n), hence $a(a(f(n)))=a(n)$. ∎
4.5 Wythoff endpoints
The endpoints of the Beatty interval I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}{-}1)] coincide with the parameters of the Wythoff complementary pair.
Theorem 18.
Let $\alpha=\sqrt{2}$.
  1. 1.
    At $d=\alpha/2$, $a(n)=\lfloor n/\alpha+\alpha/2\rfloor$ and the complement $\bar{a}(n)=\lfloor n\alpha+\alpha/2\rfloor$ form a Beatty partition of $\mathbb{Z}_{\geq 1}$.
  2. 2.
    At $d=2(\alpha{-}1)$, $a(n)=\lfloor cn+2(\alpha{-}1)\rfloor$ and $\bar{a}(n)=n+a(n)$ form a complementary pair in the $\alpha$-Wythoff game.
The lower endpoint gives the standard Beatty partition via Rayleigh’s theorem (since $1/\alpha+1/(1+\alpha)=1$). The upper endpoint is the $\alpha$-Wythoff pair studied by Fraenkel [4].
Remark 19.
At $d=\alpha/2$, the canonical solution satisfies both the strong identity $a(f(n))=n$ and the Beatty complementation. This double constraint pins down the shift uniquely. At the upper endpoint $d=2(\alpha{-}1)$, only the triple-nested identity holds, and the Beatty complementation is replaced by the Wythoff partition.
5 Internal structure
Throughout this section we consider the case $r=2$ and write $a_{\mathrm{B}}$ for the canonical Beatty solution (2) and $a_{\mathrm{G}}$ for the greedy solution.
5.1 Sturmian words
The difference sequence $w_{\mathrm{B}}(n)=a_{\mathrm{B}}(n{+}1)-a_{\mathrm{B}}(n)\in\{0,1\}$ is a Sturmian word of slope $c=1/\!\sqrt{2}$ and intercept $d{-}(1{-}c)=3c{-}1=(3-\sqrt{2})/\sqrt{2}$, namely $w_{\mathrm{B}}(n)=\lfloor c(n{+}1)+d\rfloor-\lfloor cn+d\rfloor$. The word is balanced, and every pair of factors of the same length differ by at most one in the number of $1$’s.
The multiplicity $m(n)=\#\{k:a_{\mathrm{B}}(k)=n\}$ takes values in $\{1,2\}$, and $m(n)=2$ iff $w_{\mathrm{B}}(n{-}1)=0$, i.e., $\{cn+d\}<1{-}c$. This characterises the positions of repeated values purely in terms of the Sturmian rotation.
A morphism on $\{0,1\}^{*}$ generating this Sturmian word (via the standard Rauzy construction [7]) is $\sigma\colon 0\mapsto 01,\;1\mapsto 011$, with $\sigma^{k}(0)\to w_{\mathrm{B}}$ as $k\to\infty$. The substitution is associated with the continued fraction $0;1,2,2,2,\ldots]$ via the classical correspondence between Sturmian morphisms and Ostrowski numeration [7, Chapter 6].
5.2 Bifurcations
The two solutions $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$ first diverge at $n=12$ (Section 1), and their difference words diverge one step earlier, with $w_{\mathrm{G}}(11)=0$ and $w_{\mathrm{B}}(11)=1$. This is the smallest instance of a broader phenomenon. For $r=2$, the greedy $a_{\mathrm{G}}$ and the Beatty $a_{\mathrm{B}}$ turn out to be only two members of a large family of monotone solutions, organised along a binary branching structure.
More precisely, a companion paper [3] proves that the set $\mathcal{S}$ of all nondecreasing solutions of (1) with $r=2$ admits a canonical bijection with the Cantor space $\{0,1\}^{\mathbb{N}}$, in which the greedy corresponds to $0^{\infty}$, the “lazy” (anti-greedy) solution to $1^{\infty}$, and the ratio $a_{\omega}(n)/n$ converges to $1/\!\sqrt{2}$ with a log-periodic residual for every $\omega$. The Beatty solution studied here lies outside this tree, and is characterised by its Sturmian difference word rather than by any binary branching pattern. A filtering theorem then selects, among the continuum of tree solutions, the countably many that are $2$-regular (exactly the eventually constant directing words). We refer the reader to [3] for statements and proofs.
6 The defect at the right endpoint
6.1 Definition
For $d=2(\sqrt{2}{-}1)$ (the right endpoint of $I$), the triple-nested identity $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n)$ holds by Theorem 3. The strong identity $a_{\mathrm{B}}(f(n))=n$ however fails at certain positions. Define the defect
$\delta(n)=a_{\mathrm{B}}(f(n))-n\in\{0,1\}$
and the defect set $\mathcal{D}=\{n\geq 1:\delta(n)=1\}$. The first elements of $\mathcal{D}$ are
[…]
Write $\theta_{n}=\{cn+d\}\in[0,1)$, so that $a_{\mathrm{B}}(n)=cn+d-\theta_{n}$.
6.2 The absorption theorem
The central result of this section is that the defect positions are characterised by a simple interval condition on the Sturmian address $\theta_{n-1}$, and in particular every defect position is a repeat of $a_{\mathrm{B}}$.
Theorem 20.
For $d=2(\sqrt{2}{-}1)$ and $n\geq 2$:
$\delta(n)=1\iff\theta_{n-1}\in[1-c,\,1/2].$
Consequently:
  1. (a)
    Every $n\in\mathcal{D}$ is a repeat of $a_{\mathrm{B}}$, i.e. $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$.
  2. (b)
    The set $\mathcal{D}$ has natural density
    $\rho(\mathcal{D})=c-\tfrac{1}{2}=\frac{\sqrt{2}-1}{2}\approx 0.2071.$
Proof.
Using $a_{\mathrm{B}}(n)=cn+d-\theta_{n}$ and $a_{\mathrm{B}}(n{-}1)=c(n{-}1)+d-\theta_{n-1}$:
$f(n)=a_{\mathrm{B}}(n)+a_{\mathrm{B}}(n{-}1)=c(2n{-}1)+2d-\theta_{n}-\theta_{n-1}.$
Since $c^{2}=1/2$:
$cf(n)+d=n-\tfrac{1}{2}+d(2c{+}1)-c(\theta_{n}+\theta_{n-1}).$
At $d=2(\sqrt{2}{-}1)=4c-2$, $d(2c{+}1)=(4c-2)(2c{+}1)=8c^{2}+2c-2=2$, so
$cf(n)+d=n+\tfrac{3}{2}-c(\theta_{n}+\theta_{n-1}).$ (19)
Applying the floor, $a_{\mathrm{B}}(f(n))=n+1$ precisely when $n{+}1\leq cf(n)+d<n+2$, which using (19) rearranges to $\tfrac{1}{2}\leq-c(\theta_{n}+\theta_{n-1})<\tfrac{1}{2}$, equivalently $\theta_{n}+\theta_{n-1}\leq 1/(2c)=c$ (the lower bound is automatic).
Now $\theta_{n}=\{\theta_{n-1}+c\}$. Since $c<1$, two cases arise.
Case $\theta_{n-1}<1-c$. Then $\theta_{n}=\theta_{n-1}+c$, and $\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c\leq c$ forces $\theta_{n-1}=0$, which never happens since $d$ is irrational.
Case $\theta_{n-1}\geq 1-c$. Then $\theta_{n}=\theta_{n-1}+c-1$, and $\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c-1\leq c$ becomes $\theta_{n-1}\leq 1/2$, yielding $\theta_{n-1}\in[1-c,1/2]$.
This interval is non-empty since 1/2>1c1/2>1-c, with length $1/2-(1-c)=c-1/2=(\sqrt{2}{-}1)/2$. The characterisation is established.
For (a), in the effective case $\theta_{n}=\theta_{n-1}+c-1\in[0,c-1/2]$. Since c1/2<1cc-1/2<1-c, we have $\theta_{n}<1-c$, so $\theta_{n}+c<1$ and $a_{\mathrm{B}}(n{+}1)-a_{\mathrm{B}}(n)=\lfloor\theta_{n}+c\rfloor=0$.
For (b), the sequence $(\theta_{n})_{n\geq 0}$ is the orbit of the irrational rotation $\theta\mapsto\{\theta+c\}$, uniformly equidistributed in $[0,1)$ by Weyl’s theorem. The set of $n$ with $\theta_{n-1}\in[1-c,1/2]$ therefore has natural density equal to the length of that interval. ∎
The absorption mechanism is now transparent: at a defect position, the Sturmian address $\theta_{n}$ lies in the low interval [0,c1/2][0,c-1/2], forcing $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$. The identity (5) then survives because the extra +1+1 in $a_{\mathrm{B}}(f(n))$ is absorbed: $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$.
6.3 The gap sequence
The characterisation $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in[1-c,1/2]\}$ expresses $\mathcal{D}$ as the return-time set of the rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2]. By the three-gap theorem (Świerczkowski [10]), the gaps take at most three values. Numerical experiments confirm these three values are $3$, 44, and $7$, with empirical frequencies
$\sqrt{2}-1,\quad 3-2\sqrt{2},\quad\sqrt{2}-1,$
and average gap $3(\sqrt{2}{-}1)+4(3{-}2\sqrt{2})+7(\sqrt{2}{-}1)=2(\sqrt{2}{+}1)$. The corresponding density $1/(2(\sqrt{2}{+}1))=(\sqrt{2}{-}1)/2$ matches Theorem 20(b).
The letter $4$ is always preceded and followed by $3$: every $4$ appears in a block $3,4,3$ (verified for $k\leq 10^{5}$).
Conjecture 21.
The gap sequence of $\mathcal{D}$ on the alphabet $\{3,4,7\}$ is a coding of the fixed point of a primitive substitution; equivalently, $\mathcal{D}$ is a morphic set.
6.4 A two-letter coarsening
Grouping the three gap values into two classes makes the structure more transparent. Encode $3,4$ as the letter $S$ (short, local gap) and $7$ as the letter $L$ (long gap). The resulting two-letter sequence has frequencies $1-(\sqrt{2}{-}1)=2-\sqrt{2}$ for $S$ and $\sqrt{2}-1$ for $L$. Both frequencies lie in $\mathbb{Q}(\sqrt{2})$, consistent with the fact that $a_{\mathrm{B}}$ is a Sturmian rotation of slope $c=1/\!\sqrt{2}$. Empirically the $S/L$ sequence appears to be Sturmian with slope $\sqrt{2}-1$, but a rigorous identification is left as an open problem.
7 Further remarks
The next remark records the existence of a large family of other nondecreasing solutions for $r=2$.
Remark 22.
The implicit nature of (1) allows multiple self-consistent continuations. For $r=2$, the resulting tree of nondecreasing solutions is a complete binary tree, and a continuum of infinite paths coexist with the greedy and the Beatty. This tree structure, together with an algebraic filter that isolates the $2$-regular paths, is developed in a separate paper [3].
8 Open problems
  1. 1.
    Is the affine solution (7) the unique continuous monotone solution of the continuous equation (6)? By Theorem 4, any such solution is a homeomorphism satisfying $\varphi^{-1}(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$.
  2. 2.
    Among nondecreasing solutions of (1) for $r\geq 3$, classify all monotone solutions. For $r=2$ the tree structure is described in [3]; the general case is open.
  3. 3.
    In the Ostrowski numeration system for $c=[0;1,2,2,2,\ldots]$ (see [9]), with convergents $p_{k}/q_{k}$ and digits $0\leq d_{k}\leq a_{k+1}$, verify that $a_{\mathrm{B}}(n)=\sum d_{k}p_{k}$ whenever $n=\sum d_{k}q_{k}$. This digit-swap property has been confirmed numerically for $n\leq 10^{6}$ and would give a carry-free interpretation of the identity $a_{\mathrm{B}}(a_{\mathrm{B}}(n)+a_{\mathrm{B}}(n{-}1))=n$.
  4. 4.
    Prove Conjecture 21 and identify the substitution (or the primitive system) governing the defect set $\mathcal{D}$ at the right endpoint of the Beatty interval, for $r=2$ and for each $r\geq 3$.