### 21 Apr 2026 v3

# Beatty solutions of almost Golomb equations

Abstract
The almost Golomb equation of order $r$ is the implicit functional equation
$a\Bigl(\sum_{j=0}^{r-1}a(n{-}j)\Bigr)=n$
for nondecreasing sequences of positive integers with $a(1)=1$. Its earliest solution, the almost Golomb sequence of order $r$, is $r$-regular in the sense of Allouche and Shallit and has oscillating ratio $a(n)/n$. We prove that for every $r\geq 2$ that is not an even perfect square, the equation admits a second monotone solution given by an inhomogeneous Beatty sequence of slope $1/\!\sqrt{r}$.
Composing the equation with $a$ leads to a triple-nested identity which admits a continuous one-parameter family of inhomogeneous Beatty solutions, parametrised by a shift $d$ ranging over an explicit interval. We determine these intervals sharply for $r=2$ and $r=3$, each proved by a local regime analysis combined with equidistribution of an irrational orbit. The endpoints of these intervals sit naturally inside the Pell–Ostrowski framework of Fokkink, and the defect set at the upper endpoint for $r=2$ is characterised as the return-time set of an irrational rotation to an explicit interval.
2020 Mathematics Subject Classification. Primary 11B83. Secondary 11B37, 39B12, 91A46.
Keywords. Beatty sequence, almost Golomb sequence, functional equation, Sturmian word, Pell tower, Ostrowski numeration.
1 Introduction
1.1 The almost Golomb equation
Golomb’s sequence [6] is the unique nondecreasing sequence of positive integers in which $n$ appears exactly $a(n)$ times. It satisfies $a(\sum_{k=1}^{n}a(k))=n$ and grows like $a(n)\sim\varphi^{2-\varphi}n^{\varphi-1}$, where $\varphi=(1+\sqrt{5})/2$ is the golden ratio.
Replacing the cumulative sum by a sliding window of size $r$ leads to the Golomb equation of order $r$
$a\Bigl(\sum_{j=0}^{r-1}a(n{-}j)\Bigr)=n,\qquad n\geq 1,$ (1)
with $a(k)=0$ for $k<1$. The earliest nondecreasing solution, called the almost Golomb sequence of order $r$ in [2], is obtained inductively: $a(n)$ is the smallest integer $\geq a(n{-}1)$ consistent with (1). This finite-memory truncation changes the nature of the sequence. The smooth power law gives way to oscillatory linear growth, and the greedy solution is $r$-regular in the sense of Allouche and Shallit [1] for every $r\geq 2$ [2]. Writing $f(n)=\sum_{j=0}^{r-1}a(n{-}j)$, equation (1) reads $a\circ f=\mathrm{id}$.
Equation (1) is implicit. It constrains $a$ at position $f(n)$, which depends on $a$ itself and typically lies far beyond $n$. This self-referential addressing creates room for multiple solutions beyond the greedy one.
1.2 The greedy solution
The greedy (earliest) solution of (1) is studied in detail in the companion paper [2]. It is unique, $r$-regular, and has explicit $r$-adic denesting formulas. For $r=2$ (A394217), it begins
[…]
and satisfies the dyadic recurrences a(2n)=a(n)+a(n+1)1a(2n)=a(n)+a(n{+}1)-1 and a(2n+1)=a(n)+a(n+1)a(2n{+}1)=a(n)+a(n{+}1) [2]. Its ratio $a(n)/n$ oscillates between $2/3$ and $3/4$ and does not converge [2].
1.3 A second solution
In this paper we show that, once the greedy constraint is removed, the same implicit equation admits a second monotone solution of a different nature. The case $r=2$ already illustrates the phenomenon.
Theorem 1.
The inhomogeneous Beatty sequence
$a(n)=\Bigl\lfloor\frac{n+1}{\sqrt{2}}\Bigr\rfloor,\qquad n\geq 1,$ (2)
with $a(k)=0$ for $k<1$, satisfies the Golomb equation of order $2$:
$a\bigl(a(n)+a(n{-}1)\bigr)=n\qquad\text{for all }n\geq 1.$ (3)
This Beatty solution begins
[…]
The two solutions agree on $[1,11]$ and diverge at $n=12$, where the greedy repeats the value $8$ (run of length $2$), while the Beatty moves to $9$ (run of length $1$). This is possible because the constraint at $n=12$ applies at position $f(12)=a(12)+a(11)$, which is $16$ for the greedy and $17$ for the Beatty. Both sequences thus satisfy the equation at $n=12$, but with the constraint applied at different positions, allowing divergence of the two solutions from that point on.
In the rest of the paper, $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$ denote the greedy and the canonical Beatty solutions for the case $r=2$. Otherwise $a$ is generic and is made precise by the context. The table below compares the two solutions and their constrained positions $S_{n}=a(n)+a(n{-}1)$ near the first divergence point.
$n$ $a_{\mathrm{G}}(n)$ $a_{\mathrm{B}}(n)$ $S^{\mathrm{G}}_{n}$ $S^{\mathrm{B}}_{n}$ $a_{\mathrm{G}}(S^{\mathrm{G}}_{n})$ $a_{\mathrm{B}}(S^{\mathrm{B}}_{n})$
10 7 7 14 14 10 10
11 8 8 15 15 11 11
12 8 9 16 17 12 12
13 9 9 17 18 13 13
14 10 10 19 19 14 14
Greedy Beatty
Rule local, minimal global, irrational slope
Structure 2-regular, dyadic Sturmian, Ostrowski
$a(n)/n$ oscillates ($2/3\leftrightarrow 3/4$) converges to $1/\!\sqrt{2}$
Multiplicity 22-automatic Sturmian
The residual $a_{\mathrm{B}}(n)-n/\!\sqrt{2}$ is bounded and equals $\{(n{+}1)/\!\sqrt{2}\}\in(-1,0]$, whereas $a_{\mathrm{G}}(n)-n/\!\sqrt{2}$ grows without bound because $a_{\mathrm{G}}(n)/n$ oscillates in $[2/3,3/4]$ rather than converging to $1/\!\sqrt{2}$.
Theorem 1 is the case $r=2$ of the following general result.
Theorem 2.
Let $r\geq 2$ and define the order-$r$ Beatty sequence
$a(n)=\Bigl\lfloor\frac{n}{\sqrt{r}}+\frac{\sqrt{r}}{2}\Bigr\rfloor.$ (4)
Then $a\bigl(\sum_{j=0}^{r-1}a(n{-}j)\bigr)=n$ for all $n\geq r$ if and only if $r$ is not an even perfect square.
When $r=s^{2}$ with $s$ even, the left-hand side equals $n+1$ instead of $n$. When $s$ is odd the identity holds as in the non-square case (Theorem 13).
The proof combines three ingredients. The continuous functional equation identifies the unique affine candidate (Section 2). Floor discretisation then reduces the identity to a finite counting inequality on fractional parts of multiples of $1/\!\sqrt{r}$ (Proposition 9). A parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$ establishes this inequality for all non-square $r$ (Section 3).
1.4 The triple-nested family
Applying $a$ to both sides of (3) gives $a(a(f(n)))=a(n)$, the triple-nested equation
$a\bigl(a(a(n)+a(n{-}1))\bigr)=a(n).$ (5)
Write the inhomogeneous Beatty sequence of slope $c=1/\!\sqrt{2}$ and shift $d$ as
$a(n)=\lfloor cn+d\rfloor.$
The canonical solution (2) corresponds to $d=\sqrt{2}/2$, which is the unique shift for which the strong identity (3) holds. The weaker triple-nested identity (5), by contrast, admits a continuous family of shifts. The paper determines this interval sharply for $r=2$, and the analogous interval for $r=3$.
Theorem 3.
Let $c=1/\!\sqrt{2}$ and I=[2/2, 2(21)]I=[$\sqrt${2}/2,\;2($\sqrt${2}-1)]. The sequence defined by $a(n)=\lfloor cn+d\rfloor$ satisfies (5) for all $n\geq 2$ if and only if $d\in I$.
The analogous theorem for $r=3$, stated and proved in Section 4.6, gives
$a(a(a(n)+a(n{-}1)+a(n{-}2)))=a(n)\quad\text{for all }n\geq 3\iff d\in I_{3}=\Bigl[\frac{\sqrt{3}+3}{6},\;\frac{5\sqrt{3}-3}{6}\Bigr].$
Both intervals are proved sharp by the same method, namely a local analysis of the four combinatorial regimes induced by the Sturmian rotation combined with equidistribution of the irrational orbit. Convergents of $c$ play no role in the proof.
The right endpoint $d=2(\sqrt{2}{-}1)$ for $r=2$ admits a simple algebraic form. At this boundary the strong identity fails on a sparse defect set. We prove (Theorem 24) that this defect set is the return-time set of the Sturmian rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1{-}c,1/2], which explains every failure by absorption into a repeat of $a_{\mathrm{B}}$. The three-distance theorem then yields three gap values $\{3,4,7\}$. Numerical evidence suggests that the gap sequence is generated by an explicit primitive substitution with Perron eigenvalue $1+\sqrt{2}$, formulated as Conjecture 26. A full proof via Rauzy induction on three letters is left for future work.
The equation, the interval $I$, and the defect structure all sit naturally inside the Pell–Ostrowski framework of Fokkink [4], which generalises the Wythoff array of Conway and Ryba to the recurrence $X_{n+1}=2X_{n}+X_{n-1}$. We make this connection explicit in Section 4.5, and note that the family $r=s^{2}+1$ for $s\geq 1$ corresponds to Fokkink’s Ostrowski towers of parameter $2s$.
1.5 Outline
Sections 23 prove the existence results. Section 2 establishes the exact continuous solution. Section 3 reduces the discrete identity to a counting inequality via sawtooth analysis, then closes it by a parity argument for all $r$ that are not even perfect squares. Section 4 is the rigorous core of the paper. It proves the Beatty intervals for the triple-nested equation at $r=2$ (Theorem 3) and at $r=3$ (Theorem 21), and describes their Pell–Ostrowski interpretation. The remaining sections discuss structural phenomena specific to $r=2$. Section 5 discusses the Sturmian structure and bifurcation behaviour. Section 6 proves that the defect set at the upper endpoint of the Beatty interval is the return-time set of a Sturmian rotation to an explicit sub-interval (Theorem 24), computes its density and the three gap values, and formulates as a conjecture the primitive substitution on $\{3,4,7\}$ whose Perron eigenvalue is the silver ratio $1+\sqrt{2}$.
2 The continuous equation
In the discrete equation (1), $f(n)=\sum a(n{-}j)$ denotes the sliding-window sum and the identity reads $a\circ f=\mathrm{id}$. The continuous analogue replaces $a$ by a continuous nondecreasing function $\varphi$:
$\varphi\Bigl(\sum_{j=0}^{r-1}\varphi(x{-}j)\Bigr)=x.$ (6)
A natural candidate is the affine function
$\varphi(x)=\frac{x}{\sqrt{r}}+\frac{\sqrt{r}-1}{2}.$ (7)
Theorem 4.
The function (7) is the unique affine solution of (6). Moreover, every continuous nondecreasing solution is a homeomorphism of $\mathbb{R}$ satisfying the inverse equation
$\varphi^{-1}(x)=\sum_{j=0}^{r-1}\varphi(x{-}j).$ (8)
Proof.
Suppose first that $\varphi(x)=cx+d$ is affine. Then
$\sum_{j=0}^{r-1}\varphi(x{-}j)=r(cx+d)-\frac{cr(r{-}1)}{2}$
and then
$\varphi\Bigl(\sum\varphi(x{-}j)\Bigr)=rc^{2}x+\Bigl(rcd-\frac{rc^{2}(r{-}1)}{2}+d\Bigr).$
This equals $x$ if and only if $rc^{2}=1$ and $d(\sqrt{r}+1)=(r{-}1)/2$. The unique positive solution is $c=1/\!\sqrt{r}$, $d=(\sqrt{r}-1)/2$.
We now show that every continuous nondecreasing solution is a homeomorphism satisfying the inverse equation. Let $\varphi$ be any such solution and set $g(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$. Then $g$ is continuous and nondecreasing, and $\varphi(g(x))=x$ for all $x$, so $\varphi$ is surjective and $g$ is injective. Since $g$ is continuous, nondecreasing, and injective, it is strictly increasing. If $\varphi$ had a flat interval $[u,v]$ with $\varphi\equiv\lambda$ there, then $g(\lambda)\in[u,v]$ but $g(x)<u$ for $x<\lambda$ and $g(x)>v$ for $x>\lambda$, contradicting continuity of $g$ at $\lambda$. So $\varphi$ is strictly increasing, hence a homeomorphism, and $g=\varphi^{-1}$. ∎
Whether the inverse equation (8) already forces $\varphi$ to be affine remains open (see Section 7).
3 The general identity
Throughout this section we fix $r\geq 2$ and write $a$ for the order-$r$ Beatty sequence
$a(n)=\Bigl\lfloor\frac{n}{\alpha}+\frac{\alpha}{2}\Bigr\rfloor,$ (9)
where $\alpha=\sqrt{r}$ and $c=1/\alpha$. Set
$S(n)=\sum_{j=0}^{r-1}a(n{-}j).$
The goal of the section is the following, announced as Theorem 2 in the introduction.
Theorem 5.
For every $r\geq 2$ that is not an even perfect square, $a(S(n))=n$ for all $n\geq r$. When $r=s^{2}$ with $s$ even, $a(S(n))=n+1$ for all $n\geq r$. When $r=s^{2}$ with $s$ odd, the identity $a(S(n))=n$ holds as in the non-square case.
The proof occupies Sections 3.1 through 3.4. It proceeds in three steps. Section 3.1 reduces the identity $a(S(n))=n$ to a uniform bound on the sawtooth function $\Phi$ (Proposition 6). Sections 3.23.3 reduce that bound to a counting inequality on fractional parts of multiples of $c$, and close the inequality by a parity argument on $t=r-\lfloor\sqrt{r}\rfloor^{2}$. Section 3.4 treats the perfect-square cases separately via Hermite’s identity.
3.1 Reduction to a bound on $\Phi$
By Theorem 4, the continuous identity $\varphi(\sum\varphi(x{-}j))=x$ is exact. The only obstruction to the discrete identity $a(S(n))=n$ is the rounding error introduced by the floor. Define the sawtooth function
$\Phi(n):=\sum_{j=0}^{r-1}\Bigl\{\frac{n{-}j}{\alpha}+\frac{\alpha}{2}\Bigr\}.$
Proposition 6.
For all $n\geq r$,
$a(S(n))=n\quad\iff\quad\frac{r-\alpha}{2}<\Phi(n)\leq\frac{r+\alpha}{2}.$
Proof.
Write $x_{j}=(n{-}j)/\alpha+\alpha/2$, so that $a(n{-}j)=x_{j}-\{x_{j}\}$. Summing over $j=0,\ldots,r{-}1$:
$S(n)=\sum x_{j}-\Phi(n).$
Since $\alpha^{2}=r$ we get $\sum x_{j}=(n+1/2)\alpha$, hence $S(n)=(n+1/2)\alpha-\Phi(n)$. Applying $a$:
$a(S(n))=\Bigl\lfloor n+\frac{1+\alpha}{2}-\frac{\Phi(n)}{\alpha}\Bigr\rfloor.$
This equals $n$ iff $0\leq(1{+}\alpha)/2-\Phi(n)/\alpha<1$, which rearranges to $(r{-}\alpha)/2<\Phi(n)\leq(r{+}\alpha)/2$. ∎
3.2 Counting reformulation
Write $s=\lfloor\alpha\rfloor$, $\beta=\alpha-s\in(0,1)$, and let
$N(r)=\#\bigl\{1\leq j\leq r{-}1:\{j/\alpha\}<\beta\bigr\}.$ (10)
Lemma 7.
The pairing $j\leftrightarrow r{-}j$ gives
$2\sum_{j=1}^{r-1}\lfloor jc\rfloor=(r{-}1)(s{-}1)+N(r),$ (11)
and $N(r)<r\beta+s-1$ is equivalent to
$\sum_{j=1}^{r-1}\{jc\}>\frac{r-\alpha}{2}.$ (12)
Proof.
Since $jc+(r{-}j)c=\alpha\notin\mathbb{Z}$, $\lfloor jc\rfloor+\lfloor(r{-}j)c\rfloor=s$ when $\{jc\}<\beta$ and =s1=s{-}1 otherwise, giving (11). Writing $\sum\{jc\}=(r{-}1)\alpha/2-[(r{-}1)(s{-}1)+N(r)]/2$, the equivalence with (12) is algebraic. ∎
The local maximum of the sawtooth just before the jump at $\theta=\{j_{0}c\}$, where $j_{0}\in\{1,\ldots,r{-}1\}$ indexes the jump points, is
$M(j_{0})=r-j_{0}+\sum_{k=1}^{j_{0}}\{kc\}-\sum_{m=1}^{r-1-j_{0}}\{mc\},$ (13)
and the local minimum just after the same jump is $M(j_{0})-1$ (using $\{-mc\}=1-\{mc\}$ for $m\geq 1$).
The supremum and infimum of the sawtooth $\Phi$ enjoy a palindromic symmetry.
Lemma 8.
$\sup\Phi+\inf\Phi=r$.
Proof.
Set $\Delta_{k}=\{kc\}+\{(r{-}k)c\}-1$ for $k=1,\ldots,r{-}1$ and $P(j)=\sum_{k=1}^{j}\Delta_{k}$, with $P(0)=0$. By (13), $M(j_{0})=r-\Sigma+P(j_{0})$ where $\Sigma=\sum_{j=1}^{r-1}\{jc\}$. Since $\{kc\}+\{(r{-}k)c\}$ is symmetric in $k$ and rkr{-}k, $\Delta_{k}=\Delta_{r-k}$, so the step sequence is palindromic and P(j)+P(r1j)=P(r1)P(j)+P(r{-}1{-}j)=P(r{-}1) for all $j$. Therefore $\max P+\min P=P(r{-}1)$. Then
$\sup\Phi+\inf\Phi$ $=(r{-}\Sigma{+}\max P)+(r{-}1{-}\Sigma{+}\min P)$
$=2r{-}1{-}2\Sigma+P(r{-}1).$
Substituting $P(r{-}1)=(r{-}1)\beta-N(r)$ and $2\Sigma=(r{-}1)(\beta{+}1)-N(r)$ (from (11)) gives $2r{-}1{-}2\Sigma+P(r{-}1)=r$. ∎
Proposition 9.
For non-square $r$, $N(r)<r\beta+s-1$ implies $a(S(n))=n$ for all $n\geq r$.
Proof.
By Proposition 6, the identity $a(S(n))=n$ for all $n\geq r$ is equivalent to $(r{-}\alpha)/2<\Phi(n)\leq(r{+}\alpha)/2$ for all such $n$. Since $\theta=\{n/\alpha+\alpha/2\}$ is equidistributed in $[0,1)$ by Weyl’s theorem (using that $c=1/\alpha$ is irrational), and $\Phi(n)$ depends on $n$ only through $\theta$, the bound extends to the sawtooth $\Phi\colon[0,1)\to\mathbb{R}$ viewed as a function of the orbit parameter $\theta$. By Lemma 8, $\sup\Phi+\inf\Phi=r$, so the two bounds are equivalent: it suffices to show $\sup\Phi<(r{+}\alpha)/2$ (equality excluded by irrationality of $c$).
By (13), $M(j_{0})=r-\Sigma+P(j_{0})$ where $\Sigma=\sum\{jc\}$ and $P(j)=\sum_{k=1}^{j}\Delta_{k}$ as in Lemma 8. So $\sup\Phi=r-\Sigma+\max P$, and the required bound becomes $\max P<\Sigma-(r{-}\alpha)/2$.
Since each step $\Delta_{k}$ lies in $\{\beta{-}1,\,\beta\}$, the partial sum $P(j)=j\beta-C(j)$ where $C(j)=\#\{1\leq k\leq j:\{kc\}<\beta\}$. The orbit $(\{kc\})_{k\geq 1}$ is a Sturmian rotation, and the bounded remainder property of irrational rotations gives $|C(j)-j\beta|<1$ for all $j$ (this is the balance property of Sturmian words, see [7, Chapter 2]). Hence $|P(j)|<1$ for every $j$. By Lemma 8, $\max P+\min P=P(r{-}1)$, so $\max P-\min P<2$.
Now $\sup\Phi<(r{+}\alpha)/2$ is equivalent to $\max P-\min P<\alpha{-}1$ (substituting $\Sigma-(r{-}\alpha)/2=[\alpha{-}1{-}P(r{-}1)]/2$ and using $\max P-\min P=2\max P-P(r{-}1)$). For non-square $r\geq 10$ one has $\alpha{-}1>2$, so the bound $\max P-\min P<2$ suffices. For $3\leq r\leq 8$ non-square, the bound is checked by direct evaluation of the relevant quantities, as recorded in Table 1. For $r=2$, $\sup\Phi=(r{+}\alpha)/2$ exactly, but $c\notin\mathbb{Q}$ ensures the supremum is not attained. ∎
The quantities entering the counting inequality $N(r)<r\beta+s-1$ for the small cases $r\in\{2,3,5,6,7,8\}$ are collected below, together with the resulting extrema $\sup\Phi$ and $\inf\Phi$. The margin column $r\beta+s-1-N(r)$ is positive in every case.
$r$ $s$ $\beta$ $N(r)$ $r\beta+s-1$ margin $\sup\Phi$ $\inf\Phi$
22 11 $0.4142$ 0 $0.8284$ $0.8284$ $1.7071$ $0.2929$
33 11 $0.7321$ 22 $2.1962$ $0.1962$ $2.2679$ $0.7321$
55 22 $0.2361$ 0 $2.1803$ $2.1803$ $3.4721$ $1.5279$
66 22 $0.4495$ 33 $3.6969$ $0.6969$ $3.8763$ $2.1237$
77 22 $0.6458$ 44 $5.5203$ $1.5203$ $4.3542$ $2.6458$
88 22 $0.8284$ 77 $7.6274$ $0.6274$ $5.1005$ $2.8995$
Table 1: Direct verification of the counting inequality $N(r)<r\beta+s-1$ for small non-square $r$. The margin column is $r\beta+s-1-N(r)$, positive in every case. The extrema satisfy $\sup\Phi+\inf\Phi=r$ (Lemma 8) and both lie in the required interval $((r-\alpha)/2,(r+\alpha)/2]$. The tightest case is $r=3$ with margin $\approx 0.196$.
Remark 10.
The value $N(r)=0$ observed at $r=2$ and $r=5$ in Table 1 is not an anomaly. One has $N(r)=0$ precisely when $r=s^{2}+1$ for some $s\geq 1$ (sequence A002522), the values for which $\sqrt{r}$ has continued-fraction expansion $s;\overline{2s}]$ of minimal period one. In that case $t=1$ and the counting inequality $N(r)<r\beta+s-1$ reduces to $r\beta+s-1>0$, obvious for $s\geq 1$. The next such values are $r=10,\,17,\,26,\,37,\ldots$
3.3 The parity argument
Write $r=s^{2}+t$ with $1\leq t\leq 2s$, so that $\alpha=s+\beta$ with $\beta\in(0,1)$ and $\beta(2s+\beta)=t$. The key observation is that this relation allows us to trade the irrational parameter $\beta$ for the integer $t$, whose parity determines the floor $u=\lfloor s\beta\rfloor$ and thereby the structure of the counting bound.
Lemma 11.
$\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$.
Proof.
Since $\beta=t/(s+\alpha)$ and $s+\alpha\in(2s,2s{+}1)$, we get
$\frac{t}{2s{+}1}<\beta<\frac{t}{2s}.$
Multiplying by $s$:
$\frac{st}{2s{+}1}<s\beta<\frac{t}{2}.$
Since
$\frac{st}{2s{+}1}-\frac{t{-}1}{2}=\frac{2s{-}t{+}1}{2(2s{+}1)}>0$
(as $t\leq 2s$), we obtain $(t{-}1)/2<s\beta<t/2$, whence $\lfloor s\beta\rfloor=\lfloor(t{-}1)/2\rfloor$. ∎
Lemma 12.
$N(r)=(s{+}1)(t{-}1)-2\sum_{k=1}^{s}\lfloor k\beta\rfloor$.
Proof.
For $0\leq m\leq s{-}1$, the number of integers $j\in[1,r{-}1]$ with $\lfloor j/\alpha\rfloor=m$ and $\{j/\alpha\}<\beta$ is $C_{m}=t+\lfloor(m{-}s)\beta\rfloor-\lfloor m\beta\rfloor$ (since $m\alpha+\alpha\beta=m\alpha+t-s\beta$ and the endpoints are irrational). For $m=s$: $C_{s}=t{-}1{-}\lfloor s\beta\rfloor$. Using $\lfloor-m\beta\rfloor=-\lfloor m\beta\rfloor-1$ for $m\geq 1$ and summing gives the result. ∎
Since $\lfloor k\beta\rfloor$ counts integers $m\geq 1$ with $m\leq k\beta$, swapping the order of summation gives
$\sum_{k=1}^{s}\lfloor k\beta\rfloor=\sum_{m=1}^{u}\bigl(s{+}1-\lceil m/\beta\rceil\bigr),$ (14)
where $u=\lfloor s\beta\rfloor$.
We now prove the non-square part of Theorem 5.
Proof of Theorem 5, non-square case.
By Proposition 9, it suffices to show $N(r)<r\beta+s-1$. By Lemmas 11 and 12, we split by the parity of $t$.
Suppose first that t=2u+1t=2u{+}1 is odd. Then $\lfloor s\beta\rfloor=u$, and substituting (14) into Lemma 12 gives $N(r)=2\sum_{m=1}^{u}\lceil m/\beta\rceil$. Since $\beta$ is irrational, each $\lceil m/\beta\rceil<m/\beta+1$, so
$N(r)<\frac{u(u{+}1)}{\beta}+2u.$
The margin is
$r\beta+s-1-\frac{u(u{+}1)}{\beta}-2u=\frac{(1{-}\beta)^{2}(\beta^{2}{+}2\beta{+}4u{+}3)}{4\beta}>0$
(the equality is checked by substituting $r=s^{2}+t$ and $\beta(2s{+}\beta)=t=2u{+}1$, then clearing denominators).
Suppose now that $t=2u$ is even. Then $\lfloor s\beta\rfloor=u{-}1$, and the same substitution gives $N(r)=s{+}1+2\sum_{m=1}^{u-1}\lceil m/\beta\rceil$. Using $\lceil m/\beta\rceil<m/\beta+1$ again:
$N(r)<s+1+\frac{u(u{-}1)}{\beta}+2u-2.$
The margin is
$r\beta+s-1-s-1-\frac{u(u{-}1)}{\beta}-2u+2=\frac{u(1{-}\beta)^{2}}{\beta}+\frac{\beta^{3}}{4}>0$
(checked by substituting $\beta(2s{+}\beta)=t=2u$ and clearing denominators).
In both cases $N(r)<r\beta+s{-}1$ and the theorem follows from Proposition 9. ∎
3.4 Perfect squares
When $r=s^{2}$, the parameter $\beta$ vanishes and the counting argument of the previous subsection degenerates. The window sum can instead be evaluated exactly by grouping terms modulo $s$, which leads to a clean application of Hermite’s identity.
Theorem 13.
Let $r=s^{2}$ and $a(n)=\lfloor n/s+s/2\rfloor$. For all $n\geq s^{2}$:
  1. 1.
    if $s$ is odd, S(n)=sns(s1)/2S(n)=sn-s(s{-}1)/2 and $a(S(n))=n$;
  2. 2.
    if $s$ is even, S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2 and $a(S(n))=n+1$.
Proof.
Write $q=\lfloor s/2\rfloor$.
Suppose first that s=2q+1s=2q{+}1 is odd. Then $d=s/2=q+1/2\notin\mathbb{Z}$, and $a(n)=q+\lfloor(n{+}q)/s\rfloor$. For $n\geq s^{2}$ the full window is active, so
$S(n)=s^{2}q+\sum_{j=0}^{s^{2}-1}\Bigl\lfloor\frac{n-j+q}{s}\Bigr\rfloor.$
Write $j=ks+u$ with $0\leq k,u\leq s{-}1$. Then $\lfloor(n{-}ks{-}u{+}q)/s\rfloor=\lfloor(n{-}u{+}q)/s\rfloor-k$, giving
$S(n)=s^{2}q+s\sum_{u=0}^{s-1}\Bigl\lfloor\frac{n{-}u{+}q}{s}\Bigr\rfloor-s\cdot\frac{s(s{-}1)}{2}.$
As $u$ ranges over $\{0,\ldots,s{-}1\}$, the values n+qun{+}q{-}u hit each residue class modulo $s$ exactly once, so by Hermite’s identity ($\sum_{v=0}^{n-1}\lfloor x+v/n\rfloor=\lfloor nx\rfloor$ for all $x\in\mathbb{R}$)
$\sum_{u=0}^{s-1}\Bigl\lfloor\frac{n{+}q{-}u}{s}\Bigr\rfloor=n+q-s+1.$
Substituting and using q=(s1)/2q=(s{-}1)/2 yields S(n)=sns(s1)/2S(n)=sn-s(s{-}1)/2. Finally, $a(S(n))=q+\lfloor n{-}q+q/s\rfloor=q+(n{-}q)=n$, since $0\leq q/s=(s{-}1)/(2s)<1$.
Suppose now that $s=2m$ is even. Then $d=m\in\mathbb{Z}$ and $a(n)=m+\lfloor n/s\rfloor$. Applying Hermite’s identity in the same way gives S(n)=sns(s2)/2S(n)=sn-s(s{-}2)/2. Then $a(S(n))=m+\lfloor n{-}(s{-}2)/2\rfloor=n+1$, since $(s{-}2)/2=m{-}1\in\mathbb{Z}$. ∎
Corollary 14.
The canonical Beatty identity $a(S(n))=n$ holds for all odd perfect squares $r=s^{2}$ ($s=3,5,7,\ldots$) and fails systematically for all even perfect squares $r=(2m)^{2}$.
3.5 Coexistence of solutions
For each $r$ (excluding $r=(2m)^{2}$), the Golomb equation (1) of order $r$ admits at least two different monotone solutions. The greedy [2] is $r$-regular with oscillating ratio $a(n)/n$. The Beatty has $a(n)/n\to 1/\!\sqrt{r}$. Both are fixed points of the implicit equation $a\circ f=\mathrm{id}$, sustained by different self-consistent address structures.
Remark 15.
The greedy branch is governed by finite-state local structure and is therefore amenable to $r$-regular methods [2, 1]. The Beatty branch is governed by irrational rotation and the continuous functional equation (6).
4 The triple-nested equation
We now compose the Golomb equation (1) for $r=2$ with $a$, obtaining the weaker identity $a(a(f(n)))=a(n)$. This reveals a continuous family of Beatty solutions parametrised by a shift $d$ ranging over an explicit interval. The goal of this section is the following, announced as Theorem 3 in the introduction.
Theorem 16.
Let $c=1/\!\sqrt{2}$, f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1), and $a(n)=\lfloor cn+d\rfloor$. The triple-nested identity $a(a(f(n)))=a(n)$ holds for all $n\geq 2$ if and only if
$d\in I=\bigl[\sqrt{2}/2,\;2(\sqrt{2}-1)\bigr].$
The proof occupies Sections 4.14.4. Necessity is handled by two failure lemmas that locate, in each of the three residue classes of $\theta_{n}=\{cn+d\}$, a sub-interval where the identity (5) fails (Section 4.3). Orbit density, via Weyl equidistribution, then produces infinitely many failures. Sufficiency is a case analysis of the same residue classes (Section 4.4). Section 4.5 then places the interval II and its endpoints inside Fokkink’s Pell–Ostrowski framework. Section 4.6 proves the analogous result for $r=3$ by the same method.
4.1 The continuous equation
For the strong identity $\varphi(g(x))=x$ and the triple-nested $\varphi(\varphi(g(x)))=\varphi(x)$ (where $g(x)=\sum\varphi(x{-}j)$), the continuous equation gives the same unique affine solution (7) in both cases: if $\varphi(g(x))=x$, composing with $\varphi$ immediately yields $\varphi(\varphi(g(x)))=\varphi(x)$, and conversely, injectivity of $\varphi$ recovers the strong identity from the triple-nested one.
In the discrete setting, the floor breaks this equivalence. The strong identity pins down a unique shift $d=\sqrt{r}/2$ (the left endpoint of the Beatty interval), while the weaker triple-nested identity tolerates a range of shifts. The floor in the outer composition $a(\cdot)$ absorbs small errors, opening the interval $I$ to the right.
4.2 Expansion
Throughout this section, $c=1/\!\sqrt{2}$, $d\in\mathbb{R}$, and $a(n)=\lfloor cn+d\rfloor$. Write $\theta_{n}=\{cn+d\}\in[0,1)$, so that $a(n)=cn+d-\theta_{n}$, and say that index $n$ is a repeat when a(n+1)=a(n)a(n{+}1)=a(n), and a non-repeat when a(n+1)=a(n)+1a(n{+}1)=a(n){+}1. Since $a(n{+}1)-a(n)=\lfloor\theta_{n}+c\rfloor$, this amounts to $\theta_{n}<1{-}c$ (repeat) or $\theta_{n}\geq 1{-}c$ (non-repeat). Set also $e(n)=-\theta_{n}\in(-1,0]$, so that $a(n)=cn+d+e(n)$. Define f(n)=a(n)+a(n1)f(n)=a(n)+a(n{-}1). Then
f(n)=c(2n1)+2d+e(n)+e(n1).f(n)=c(2n{-}1)+2d+e(n)+e(n{-}1). (15)
Applying $a$ to $f(n)$ and using $2c^{2}=1$:
$a(f(n))=\frac{2n{-}1}{2}+d(2c{+}1)+c\bigl(e(n)+e(n{-}1)\bigr)+e(f(n)).$ (16)
A second application of $a$ yields $a(a(f(n)))=cn+R(n)$, where
$R(n)=R_{0}+c^{2}(e(n)+e(n{-}1))+c\,e(f(n))+e(a(f(n))),$ (17)
with constant part
$R_{0}=d(c+2)-\frac{c}{2}.$ (18)
The equation (5) holds iff $\lfloor cn+R(n)\rfloor=\lfloor cn+d\rfloor$ for all $n\geq 1$.
4.3 Necessity
Any solution must satisfy $a(1)=\lfloor c+d\rfloor=1$, requiring $1-c\leq d<2-c$. The failure of (5) outside $I$ is produced by a local argument on the three residue classes of $\theta_{n}\in[0,1)$ combined with density of the orbit $(\theta_{n})_{n\geq 1}$. We describe the dynamics once, and then read off each endpoint as a special case.
Expansion (16) together with $2c^{2}=1$ gives
$a(f(n))=\lfloor n-\tfrac{1}{2}+D-c\Theta_{2}(n)\rfloor,\qquad D=d(2c+1),\quad\Theta_{2}(n)=\theta_{n}+\theta_{n-1}.$ (19)
Writing $a(f(n))=n+m(n)$, the integer $m(n)$ is determined by
$m(n)=k\iff D-c\Theta_{2}(n)\in\Bigl[k+\tfrac{1}{2},\,k+\tfrac{3}{2}\Bigr).$
Two simple implications are immediate. If $m(n)\geq 1$, then $a(f(n))\geq n+1$, so survival of the identity $a(a(f(n)))=a(n)$ forces $a(n+1)=a(n)$, that is, $n$ is a repeat. Symmetrically, if $m(n)\leq-1$ then n1n-1 must be a repeat. These implications are the two basic necessary conditions
  • (N+)
    $m(n)\geq 1\ \Rightarrow\ a(n+1)=a(n)$ ($n$ is a repeat),
  • (N-)
    $m(n)\leq-1\ \Rightarrow\ a(n-1)=a(n)$ (n1n-1 is a repeat).
Taken as a global statement over all $d$, (N+) and (N-) are only necessary. If $|m(n)|\geq 2$, a single repeat is no longer enough to absorb the extra jump, and the identity fails even when $n$ or n1n-1 happens to be a repeat. The bound $|m(n)|\leq 1$ must therefore be ensured a priori. We show in Section 4.4 that this bound holds uniformly on the interval II, so that on $I$ the two conditions (N+) and (N-) become jointly sufficient as well. The necessity lemmas below exploit the two implications directly.
The three residue classes of $\theta=\theta_{n}$ are
$\mathrm{I}:\ \theta\in[0,1{-}c),\qquad\mathrm{II}:\ \theta\in[1{-}c,c),\qquad\mathrm{III}:\ \theta\in[c,1),$
which match the three branches of the decomposition in Lemma 20 below. In class I, $\theta_{n}<1-c$ means $n$ is a repeat. In classes II and III, $n$ is a non-repeat, so (N+) forbids $m(n)\geq 1$ there. In class II, furthermore, $\theta_{n-1}=\theta-c+1\in[2-2c,1)\subset[1-c,1)$, so n1n-1 is also a non-repeat, and (N-) forbids $m(n)\leq-1$ on class II.
Write $\Theta_{2}(n)$ explicitly in each class. In class III, $\theta_{n-1}=\theta-c\in[0,1-c)$, hence $\Theta_{2}(n)=2\theta-c$ and
$c\Theta_{2}(n)=2c\theta-\tfrac{1}{2}.$ (20)
In class II, $\theta_{n-1}=\theta+1-c$, hence $\Theta_{2}(n)=2\theta+1-c$ and
$c\Theta_{2}(n)=2c\theta+c-\tfrac{1}{2}.$ (21)
The orbit $(\theta_{n})_{n\geq 1}$ is the forward orbit of $\{c+d\}$ under the irrational rotation $\theta\mapsto\{\theta+c\}$. By Weyl’s equidistribution theorem, this orbit is equidistributed in $[0,1)$, so it visits every non-empty open sub-interval of $[0,1)$ at infinitely many indices $n$.
Lemma 17.
If $d<\sqrt{2}/2$, then (5) fails at infinitely many $n$.
Proof.
Compute $D$ at $d=c$: $D=c(2c+1)=2c^{2}+c=1+c$. So $d<c$ is equivalent to $D<1+c$. Work in class II, where $\theta\in[1{-}c,c)$ and $c\Theta_{2}=2c\theta+c-1/2$ by (21). As $\theta\to c^{-}$, $c\Theta_{2}\to 2c^{2}+c-1/2=c+1/2$, and so
$D-c\Theta_{2}\ \longrightarrow\ D-c-\tfrac{1}{2}.$
The hypothesis $D<1+c$ gives Dc1/2<1/2D-c-1/2<1/2, so the inequality $D-c\Theta_{2}(n)<1/2$ defines an open sub-interval of class II accumulating at $\theta=c$. Explicitly, solving $D-c\Theta_{2}<1/2$ for $\theta$ yields
$\theta>\frac{D-c}{2c},$ (22)
and the right-hand side is <c<c by hypothesis. For every $n$ with $\theta_{n}$ in this sub-interval, $m(n)\leq-1$ (and in fact m(n)=1m(n)=-1, since $D\geq c$ from the lower bound $d\geq 1-c$ keeps $D-c\Theta_{2}$ above 1/2-1/2 in this range).
Both $n$ and n1n-1 are non-repeats throughout class II, so (N-) is violated at each such $n$. Density of the orbit provides infinitely many such $n$, and (5) fails at infinitely many $n$. ∎
Lemma 18.
If $d>2(\sqrt{2}-1)$, then (5) fails at infinitely many $n$.
Proof.
Set $d_{R}=2(\sqrt{2}-1)=4c-2$. Compute $D$ at $d_{R}$:
$D(d_{R})=(4c-2)(2c+1)=8c^{2}+4c-4c-2=4-2=2,$
using $2c^{2}=1$. So $d>d_{R}$ is equivalent to $D>2$. Work in class III, where $\theta\in[c,1)$ and $c\Theta_{2}=2c\theta-1/2$ by (20). As $\theta\to c^{+}$, $c\Theta_{2}\to 2c^{2}-1/2=1/2$, and so
$D-c\Theta_{2}\ \longrightarrow\ D-\tfrac{1}{2}.$
The hypothesis $D>2$ gives D1/2>3/2D-1/2>3/2, so the inequality $D-c\Theta_{2}(n)\geq 3/2$ defines an open sub-interval of class III accumulating at $\theta=c$. Explicitly, solving $D-c\Theta_{2}\geq 3/2$ for $\theta$ yields
$\theta\leq\frac{D-1}{2c},$ (23)
and the right-hand side is >c>c by hypothesis. For every $n$ with $\theta_{n}$ in this sub-interval, $m(n)\geq 1$.
Throughout class III, $n$ is a non-repeat, so (N+) is violated. Density of the orbit supplies infinitely many such $n$, and (5) fails at infinitely many $n$. ∎
Remark 19.
The two endpoints $d=c$ and $d=d_{R}$ survive because the limiting condition would require $\theta=c$ exactly. This forces $\{cn+d\}=c$ for some integer $n$, hence $c(n-1)+d\in\mathbb{Z}$, impossible since $c$ is irrational and $d\in\mathbb{Q}(\sqrt{2})\setminus\mathbb{Q}$ at both endpoints. The orbit is dense, and irrationality excludes the single critical address at each endpoint. This dynamical argument makes the two proofs symmetric and self-contained, despite the asymmetry of $I$.
4.4 Sufficiency
The sufficiency of $d\in I$ rests on an absorption mechanism: the first nesting $a(f(n))$ may exceed $n$ by one at repeat positions, and this excess is absorbed by the outer $a$ because $a(n+1)=a(n)$ at such positions. We isolate this mechanism as a lemma and then deduce the theorem.
Lemma 20.
Let $c=1/\!\sqrt{2}$, $a(n)=\lfloor cn+d\rfloor$, and f(n)=a(n)+a(n1)f(n)=a(n)+a(n-1). For every $d\in I$ and every $n\geq 2$, $a(f(n))\in\{n,n+1\}$, with the finer dichotomy
  • if $\theta_{n}\geq 1-c$ (i.e. $n$ is a non-repeat), then $a(f(n))=n$;
  • if $\theta_{n}<1-c$ (i.e. $n$ is a repeat), then $a(f(n))\in\{n,n+1\}$.
Proof.
Write $\theta=\theta_{n}$ and D=d(2c+1)D=d(2c{+}1). We distinguish three residue classes of $\theta$.
Suppose first that $\theta\in[c,1)$. Then $f(n)=2a(n)$ and $a(f(n))=\lfloor n+\varphi\rfloor$ with $\varphi=D-2c\theta\in(D-2c,\;D-1]$. Since $d\geq c$ gives $\varphi>0$ and $d\leq 4c{-}2$ gives $\varphi\leq 1$ with equality only when d=4c2d=4c{-}2 and $\theta=c$ (requiring $c(n{+}3)\in\mathbb{Z}$, impossible), $\varphi\in(0,1)$ and $a(f(n))=n$.
Suppose next that $\theta\in[1{-}c,c)$. Then f(n)=2a(n)1f(n)=2a(n)-1 and $\varphi=D-c(2\theta{+}1)\in(0,1)$ by the same bounds on $D$, hence $a(f(n))=n$. The two sub-cases cover non-repeat positions.
Suppose finally that $\theta\in[0,1{-}c)$, so $n$ is a repeat. Then $\varphi\in(0,2)$, whence $a(f(n))\in\{n,n{+}1\}$. ∎
Proof of Theorems 3 and 16, sufficiency.
Lemma 20 gives $a(f(n))\in\{n,n+1\}$ for every $d\in I$ and $n\geq 2$, so the shift $m(n)$ introduced in Section 4.3 takes values in $\{0,1\}$. The a priori bound $|m(n)|\leq 1$ discussed there therefore holds on $I$, and the conditions (N+) and (N-) become sufficient as well as necessary. Concretely, the lemma gives $a(f(n))=n$ at non-repeat positions, so $a(a(f(n)))=a(n)$ trivially. At repeat positions $a(f(n))\in\{n,n+1\}$, and since $n$ is a repeat, a(n+1)=a(n)a(n{+}1)=a(n), so $a(a(f(n)))=a(n)$ in both subcases. The triple-nested identity therefore holds for every $n\geq 2$ and every $d\in I$. ∎
4.5 The Ostrowski–Pell framework behind the $r=2$ triple-nested family
The endpoint formulas, the return-time structure of the defect set, and the underlying irrational rotation are naturally expressed in the Ostrowski–Pell framework attached to the quadratic irrational $\alpha=1+\sqrt{2}$, organised by Fokkink [4] as a continuation of the Wythoff array of Conway and Ryba.
Shared continued-fraction backbone
The slope $c=1/\!\sqrt{2}$ has continued-fraction expansion
$c=[0;1,2,2,2,\ldots],$
whose convergents $p_{k}/q_{k}$ have denominators
$q_{k}=1,3,7,17,41,99,239,577,\ldots$
(A001333), the Pell denominators satisfying the recurrence $q_{k+1}=2q_{k}+q_{k-1}$. The characteristic equation $x^{2}=2x+1$ of this recurrence has dominant root $\alpha=1+\sqrt{2}$, the silver ratio. Fokkink’s Pell tower is built from the same recurrence as the $m=2$ instance of the family of bi-infinite sequences
$X_{n+1}=m\,X_{n}+X_{n-1},\qquad m\geq 1,$
where $m$ is a positive-integer parameter of the tower, unrelated to the shift $d$ of our Beatty ansatz. The case $m=1$ recovers the Fibonacci recurrence and the Wythoff array of Morrison and Kimberling.
Shared numeration backbone and the wall column
Every positive integer admits a unique Ostrowski representation
$n=\sum_{k\geq 1}e_{k}q_{k}$
with digits $e_{1}\in\{0,1\}$, $e_{k}\in\{0,1,2\}$ for $k\geq 2$, and $e_{k}=2\Rightarrow e_{k-1}=0$. In the tower picture, this numeration labels the rows. Fokkink distinguishes two natural Beatty sequences attached to the tower. The first column $A_{1,n}$ is the non-homogeneous Beatty sequence
$A_{1,n}=\Bigl\lfloor\frac{n\alpha}{\alpha-1}-\frac{1}{\alpha(\alpha-1)}\Bigr\rfloor=\Bigl\lfloor n\cdot(1{+}c)-\tfrac{1}{2}(2{-}\sqrt{2})\Bigr\rfloor$
of slope $\alpha/(\alpha-1)=1+c\approx 1.707$ ([4, Corollary 2.4]). The wall column $A_{m,0}$ is the non-homogeneous Beatty sequence
$A_{m,0}=\Bigl\lfloor\frac{m\alpha}{\alpha+1}\Bigr\rfloor=\Bigl\lfloor\frac{m}{\sqrt{2}}\Bigr\rfloor$ (24)
of slope $\alpha/(\alpha+1)=1/\sqrt{2}=c$ ([4, Corollary 2.5], and for $d=2$ this is A049472). Our canonical Beatty solution
$a_{\mathrm{B}}(n)=\bigl\lfloor(n{+}1)/\!\sqrt{2}\bigr\rfloor$
is a shifted version of the wall column (24), not of the first column. The difference word of $a_{\mathrm{B}}$ is Sturmian of slope $c$, and its multiplicity structure gives a Rayleigh–Fraenkel complementary partition of $\mathbb{Z}_{\geq 1}$ with slopes $1+\sqrt{2}$ and $(2+\sqrt{2})/2$ (see [5]).
Endpoint and defect interpretation
The endpoints of the interval $I$, their failure mechanisms, and the boundary defect set all read naturally in this framework. The left endpoint $d=\sqrt{2}/2$ is the unique shift for which the strong identity $a(f(n))=n$ holds, producing exactly the wall column (24) (shifted). The right endpoint $d=2(\sqrt{2}-1)=2/(1+\sqrt{2})$ is the maximal shift for which the addressing defect of the first nesting is still absorbed by the repeat structure of the underlying Sturmian difference word. Both endpoints lie in $\mathbb{Q}(\sqrt{2})\cap[0,1)$ with small Pell denominators. The failure of (5) at each endpoint is produced by orbit density near a single critical address $\theta=c$, as in Lemmas 17 and 18. The Pell-convergent identities
$cq_{k}=p_{k}+(-1)^{k}\varepsilon_{k},\qquad 2cp_{k}=q_{k}-2c(-1)^{k}\varepsilon_{k},$
with $\varepsilon_{k}=|cq_{k}-p_{k}|\to 0$, provide explicit Ostrowski witnesses of the dense orbit approaching $c$. The defect set at the upper endpoint, studied in Section 6, is the return-time set of the irrational rotation $\theta\mapsto\{\theta+c\}$ to the explicit interval [1c,1/2][1-c,1/2], and its gap alphabet $\{3,4,7\}$ consists of Pell-convergent denominators.
Generalisation to $r=s^{2}+1$
Remark 10 singled out the family $r=s^{2}+1$, $s\geq 1$, as the values of $r$ for which $N(r)=0$. This is exactly the family for which $c=1/\!\sqrt{r}$ has continued-fraction expansion $0;s,\overline{2s}]$ of minimal period one, and the corresponding Ostrowski numeration is Fokkink’s $m$-numeration with $m=2s$. The Pell tower ($s=1$) is the simplest case. Numerical exploration indicates that for every $s\geq 1$ the triple-nested identity admits a non-trivial interval of shifts. The endpoints and the associated defect structures for $s\geq 2$ are left for future work.
4.6 The triple-nested identity for $r=3$
The same mechanism extends to $r=3$. The three sliding terms produce a richer regime decomposition, and the repeat structure of $n$ and of n1n-1 no longer coincide, but the underlying logic (local regimes plus orbit density) gives a full if and only if.
Fix $r=3$, $c=1/\!\sqrt{3}$, and $a(n)=\lfloor cn+d\rfloor$. Write f(n)=a(n)+a(n1)+a(n2)f(n)=a(n)+a(n{-}1)+a(n{-}2) and $\Theta(n)=\theta_{n}+\theta_{n-1}+\theta_{n-2}$ with $\theta_{n}=\{cn+d\}$. Since $3c^{2}=1$,
$cf(n)+d=n-1+D-c\Theta(n),\qquad D=d(\sqrt{3}+1),$ (25)
and $a(f(n))=n-1+\lfloor D-c\Theta(n)\rfloor$. Setting $a(f(n))=n+m(n)$, the shift is
$m(n)=\lfloor D-c\Theta(n)\rfloor-1.$ (26)
As for $r=2$, the triple-nested identity $a(a(f(n)))=a(n)$ implies the two basic necessary conditions
  • (N+)
    $m(n)\geq 1\ \Rightarrow\ a(n+1)=a(n)$ ($n$ is a repeat),
  • (N-)
    $m(n)\leq-1\ \Rightarrow\ a(n-1)=a(n)$ (n1n-1 is a repeat),
but they are jointly equivalent to the identity only under the a priori bound $|m(n)|\leq 1$. We show below that this bound holds uniformly on $I_{3}$, so that on $I_{3}$ the necessity and the sufficiency parts can be read off from (N+) and (N-) together.
Theorem 21.
The sequence $a(n)=\lfloor cn+d\rfloor$ with $c=1/\!\sqrt{3}$ satisfies $a(a(f(n)))=a(n)$ for all $n\geq 3$ if and only if
$d\in I_{3}=\Bigl[\,\frac{\sqrt{3}+3}{6},\;\frac{5\sqrt{3}-3}{6}\,\Bigr].$ (27)
The interval is centred at $\sqrt{3}/2$ with half-width $(2\sqrt{3}-3)/6$, and both endpoints lie in $\mathbb{Q}(\sqrt{3})$.
The proof occupies this subsection. We describe the four combinatorial regimes, prove sufficiency, then prove necessity by accumulating the obstruction at $\theta=1-c$.
The four regimes
Parametrise by $t=\theta_{n-2}\in[0,1)$, so that $\theta_{n-1}=\{t+c\}$ and $\theta_{n}=\{t+2c\}$. Since $c\in(1/2,1)$ and $2c\in(1,2)$, the wraps occur at t=1ct=1-c for $\theta_{n-1}$ and at t=22ct=2-2c and t=12c(<0)t=1-2c\,(<0) for $\theta_{n}$. A direct case analysis against the threshold $\theta_{\cdot}<1-c$ that defines a repeat yields the four regimes
[…]
The sum $\Theta(n)=t+\theta_{n-1}+\theta_{n}$ is piecewise affine in $t$ and takes the values
$\Theta(t)=\begin{cases}3t+3c-1,&t\in\mathrm{A}\cup\mathrm{B},\\[2.0pt] 3t+3c-2,&t\in\mathrm{C},\\[2.0pt] 3t+3c-3,&t\in\mathrm{D}.\end{cases}$ (28)
The ranges of $\Theta$ on the four regimes are
$\Theta(\mathrm{A})$ $=[\sqrt{3}-1,5-2\sqrt{3}),$ (29)
$\Theta(\mathrm{B})$ $=[5-2\sqrt{3},2),$
$\Theta(\mathrm{C})$ $=[1,4-\sqrt{3}),$
$\Theta(\mathrm{D})$ $=[3-\sqrt{3},\;\sqrt{3}).$
Two facts drive the proof.
Fact A (lower extremum). The non-repeat positions for $n$ are exactly $\mathrm{B}\cup\mathrm{C}$, and
$\inf_{t\in\mathrm{B}\cup\mathrm{C}}\Theta(t)=1,$
attained only at the left endpoint t=1ct=1-c of regime C.
Fact B (upper extremum). The non-repeat positions for n1n-1 are exactly $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$, and
$\sup_{t\in\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}}\Theta(t)=2,$
approached only at the left limit $t\to(1-c)^{-}$ from regime B.
Both facts follow from (29) by direct inspection.
Two key threshold values for $D$
Compute $D$ at the two endpoints of $I_{3}$. At $d_{L}=(\sqrt{3}+3)/6$:
$D_{L}=\frac{\sqrt{3}+3}{6}\cdot(\sqrt{3}+1)=\frac{(\sqrt{3}+3)(\sqrt{3}+1)}{6}=\frac{3+\sqrt{3}+3\sqrt{3}+3}{6}=\frac{6+4\sqrt{3}}{6}=1+\frac{2}{\sqrt{3}}=1+2c.$
At $d_{R}=(5\sqrt{3}-3)/6$:
$D_{R}=\frac{5\sqrt{3}-3}{6}\cdot(\sqrt{3}+1)=\frac{15+5\sqrt{3}-3\sqrt{3}-3}{6}=\frac{12+2\sqrt{3}}{6}=2+\frac{1}{\sqrt{3}}=2+c.$
So $d\in I_{3}$ is equivalent to $D\in[1+2c,\,2+c]$.
Sufficiency
Assume $d\in I_{3}$, equivalently $D\in[1+2c,\,2+c]$. We first check that the shift $m(n)$ is bounded a priori, so that (N+) and (N-) together become sufficient. The sum $\Theta(n)$ ranges over $\Theta(\mathrm{A})\cup\Theta(\mathrm{B})\cup\Theta(\mathrm{C})\cup\Theta(\mathrm{D})\subset[\sqrt{3}-1,\,4-\sqrt{3})$ by (29). Multiplying by $c$ and using $c\sqrt{3}=1$,
$c\Theta(n)\in[1-c,\,4c-1),$
whence
$D-c\Theta(n)\in(D-4c+1,\,D-1+c]\subset(1+2c-4c+1,\,2+c-1+c]=(2-2c,\,1+2c].$
In particular $D-c\Theta(n)\in(0,3)$, so $\lfloor D-c\Theta(n)\rfloor\in\{0,1,2\}$ and by (26)
$m(n)\in\{-1,0,1\}\qquad\text{for every }n\geq 3,\ d\in I_{3}.$ (30)
With (30) at hand, (N+) and (N-) jointly imply the triple-nested identity on $I_{3}$. We now verify them regime by regime.
Requirement (N+) forbids $m(n)\geq 1$ at non-repeats of $n$, which by (26) is the inequality $D-c\Theta(n)\geq 2$, equivalently
$\Theta(n)\leq\frac{D-2}{c}=\sqrt{3}(D-2).$ (31)
The hypothesis $D\leq 2+c$ gives $\sqrt{3}(D-2)\leq\sqrt{3}c=1$.
For $D<2+c$ (the interior of $I_{3}$ on the right), the bound in (31) is strict, $\sqrt{3}(D-2)<1$, and Fact A gives $\Theta(n)\geq 1>\sqrt{3}(D-2)$ on $\mathrm{B}\cup\mathrm{C}$. Hence (N+) holds vacuously.
For $D=2+c$ (the right endpoint $d=d_{R}$), equality in Fact A is possible only at the single point t=1ct=1-c. Writing $d_{R}=(5c-1)/2$, the equation $\{c(n-2)+d_{R}\}=1-c$ unfolds to $c(n-1)+(5c-1)/2-1\in\mathbb{Z}$. Multiplying by $2$, this reads $(2n+3)c=2k+3$ for some integer $k$. Since $2n+3\neq 0$ and $c$ is irrational, no integer $n$ solves this equation, the threshold is never attained, and (N+) holds.
Requirement (N-) forbids $m(n)\leq-1$ at non-repeats of n1n-1, which by (26) is the inequality $D-c\Theta(n)<1$, equivalently
$\Theta(n)>\frac{D-1}{c}=\sqrt{3}(D-1).$ (32)
The hypothesis $D\geq 1+2c$ gives $\sqrt{3}(D-1)\geq 2\sqrt{3}c=2$, and by Fact B $\Theta(n)\leq 2$ on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$ with the supremum not attained. For $D>1+2c$ the bound is strict and (N-) holds vacuously. At $D=1+2c$ (the left endpoint $d=d_{L}=(1+c)/2$), the bound becomes $\Theta(n)>2$, which is impossible on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$ since the supremum $2$ is not attained. Hence (N-) holds.
Sufficiency is established for every $d\in I_{3}$.
Necessity
Right bound. Assume $d>d_{R}$, equivalently $D>2+c$. Then $\sqrt{3}(D-2)>1$, so the threshold $\sqrt{3}(D-2)$ lies strictly inside the range $\Theta(\mathrm{C})=[1,4-\sqrt{3})$, exceeding the left endpoint. The set
$U_{R}=\{t\in\mathrm{C}:\Theta(t)<\sqrt{3}(D-2)\}=\{t:1-c\leq t<(1-c)+(D-2-c)/\sqrt{3}\}$
is a non-empty open sub-interval of regime C accumulating at t=1ct=1-c. On $U_{R}$, we have $D-c\Theta(n)>2$, hence $m(n)\geq 1$, while $n$ is a non-repeat (regime C). This violates (N+). The orbit $(\theta_{n-2})_{n\geq 3}$ is dense in $[0,1)$ by Weyl’s theorem, so it visits $U_{R}$ infinitely often, and (5) fails at infinitely many $n$.
Left bound. Assume $d<d_{L}$, equivalently $D<1+2c$. Then $\sqrt{3}(D-1)<2$, so the threshold $\sqrt{3}(D-1)$ lies strictly below the supremum $2$ on $\Theta(\mathrm{B})=[5-2\sqrt{3},2)$. The set
$U_{L}=\{t\in\mathrm{B}:\Theta(t)>\sqrt{3}(D-1)\}=\{t:t_{\star}<t<1-c\},$
where $t_{\star}=\max\bigl((\sqrt{3}(D-1)+1)/3-c,\,2-3c\bigr)$, is a non-empty open sub-interval of regime B accumulating at the left limit of t=1ct=1-c. On $U_{L}$, we have $D-c\Theta(n)<1$, hence $m(n)\leq-1$, while n1n-1 is a non-repeat (regime B). This violates (N-). Density of the orbit yields infinitely many failures.
Endpoints. At $d=d_{R}$, $D=2+c$ and the threshold for $m(n)\geq 1$ is $\Theta\leq 1$. The only value in $\Theta(\mathrm{B}\cup\mathrm{C})$ satisfying this is $\Theta=1$ at t=1ct=1-c, never attained. At $d=d_{L}$, $D=1+2c$ and the threshold for $m(n)\leq-1$ is $\Theta>2$, never attained on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$. Both endpoints are therefore included in $I_{3}$.
This completes the proof of Theorem 21.
Remark 22.
The interval $I_{3}$ is symmetric around the strong-identity shift $\sqrt{3}/2$, with half-width $(2\sqrt{3}-3)/6$. Symmetry for $r=3$ contrasts with the asymmetry of $I$ for $r=2$. In both cases, the mechanism is the same. The threshold for $m(n)=+1$ and the threshold for m(n)=1m(n)=-1 are governed by extrema of $\Theta$ over the non-repeat sets of $n$ and of n1n-1 respectively, and for $r=3$ these two extrema sit symmetrically about the centre. For $r=2$, the two-term window gives only one non-repeat regime coupling both constraints, producing asymmetry.
Remark 23.
The same method applies to any non-square $r\geq 2$ with $c=1/\!\sqrt{r}$ and $f(n)=\sum_{j=0}^{r-1}a(n{-}j)$. The orbit $(\theta_{n-r+1},\dots,\theta_{n})$ partitions $[0,1)$ into finitely many regimes labelled by the joint repeat or non-repeat status of (n,n1)(n,n-1). On each regime the combined fractional-part sum $\Theta(n)$ is piecewise affine in the base parameter $t=\theta_{n-r+1}$. Local thresholds of $D-c\Theta(n)$ against the integers produce both the sufficiency bounds and the failure sets, and orbit density yields infinitely many failures outside the resulting interval $I_{r}$. We expect $I_{r}$ to be centred at $\sqrt{r}/2$, symmetric for odd $r$ and asymmetric for even $r$.
At the right endpoint $d=(5\sqrt{3}{-}3)/6$, the defect set begins
[…]
(A395252), with gap values $\{7,\,12,\,19\}$ where 19=7+1219=7{+}12. The three values correspond to consecutive convergent denominators of $1/\!\sqrt{3}=[0;1,1,2,1,2,1,2,\ldots]$, and the density of the defect set in $\mathbb{Z}_{\geq 1}$ equals $(2{-}\sqrt{3})/3$, matching the length of the rotation interval controlling the absorption. A substitution analogue of Conjecture 26 for $r=3$ is formulated as an open problem in Section 7.
5 Internal structure
Throughout this section we consider the case $r=2$ and write $a_{\mathrm{B}}$ for the canonical Beatty solution (2) and $a_{\mathrm{G}}$ for the greedy solution.
5.1 Sturmian words
The difference sequence $w_{\mathrm{B}}(n)=a_{\mathrm{B}}(n{+}1)-a_{\mathrm{B}}(n)\in\{0,1\}$ is a Sturmian word of slope $c=1/\!\sqrt{2}$ and intercept $d{-}(1{-}c)=3c{-}1=(3-\sqrt{2})/\sqrt{2}$, namely $w_{\mathrm{B}}(n)=\lfloor c(n{+}1)+d\rfloor-\lfloor cn+d\rfloor$. The word is balanced, and every pair of factors of the same length differ by at most one in the number of $1$’s.
The multiplicity $m(n)=\#\{k:a_{\mathrm{B}}(k)=n\}$ takes values in $\{1,2\}$, and $m(n)=2$ iff $w_{\mathrm{B}}(n{-}1)=0$, i.e., $\{cn+d\}<1{-}c$. This characterises the positions of repeated values purely in terms of the Sturmian rotation.
A morphism on $\{0,1\}^{*}$ generating this Sturmian word (via the standard Rauzy construction [7]) is $\sigma\colon 0\mapsto 01,\;1\mapsto 011$, with $\sigma^{k}(0)\to w_{\mathrm{B}}$ as $k\to\infty$. The substitution is associated with the continued fraction $0;1,2,2,2,\ldots]$ via the classical correspondence between Sturmian morphisms and Ostrowski numeration [7, Chapter 6].
5.2 Bifurcations
The two solutions $a_{\mathrm{G}}$ and $a_{\mathrm{B}}$ first diverge at $n=12$ (Section 1), and their difference words diverge one step earlier, with $w_{\mathrm{G}}(11)=0$ and $w_{\mathrm{B}}(11)=1$. This is the smallest instance of a broader phenomenon. For $r=2$, the greedy $a_{\mathrm{G}}$ and the Beatty $a_{\mathrm{B}}$ turn out to be only two members of a large family of monotone solutions, organised along a binary branching structure.
More precisely, a companion paper [3] proves that the set $\mathcal{S}$ of all nondecreasing solutions of (1) with $r=2$ admits a canonical bijection with the Cantor space $\{0,1\}^{\mathbb{N}}$, in which the greedy corresponds to $0^{\infty}$, the “lazy” (anti-greedy) solution to $1^{\infty}$, and the ratio $a_{\omega}(n)/n$ converges to $1/\!\sqrt{2}$ with a log-periodic residual for every $\omega$. The Beatty solution studied here lies outside this tree, and is characterised by its Sturmian difference word rather than by any binary branching pattern. A filtering theorem then selects, among the continuum of tree solutions, the countably many that are $2$-regular (exactly the eventually constant directing words). We refer the reader to [3] for statements and proofs.
6 The defect at the right endpoint
At the right endpoint $d=2(\sqrt{2}{-}1)$ of the interval $I$, the triple-nested identity holds but the strong identity $a_{\mathrm{B}}(f(n))=n$ fails at a sparse set of positions which we call the defect set $\mathcal{D}$. This section describes the structure of $\mathcal{D}$ in three steps. First we characterise $\mathcal{D}$ as the set of return times of the Sturmian rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2] (Theorem 24), which explains why every defect position is a repeat of $a_{\mathrm{B}}$, so that the extra increment in $a_{\mathrm{B}}(f(n))$ is absorbed by the outer $a$. Second, the three-distance theorem gives three possible gap values between consecutive defect positions, namely $3$, 44, and $7$. Third, the gap sequence on the alphabet $\{3,4,7\}$ is conjecturally generated by an explicit primitive substitution with Perron eigenvalue $1+\sqrt{2}$ (Conjecture 26), tying the defect structure to the Pell recurrence.
For $d=2(\sqrt{2}{-}1)$, the triple-nested identity $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n)$ holds by Theorem 3, and expansion of the defining formula shows $a_{\mathrm{B}}(f(n))\in\{n,\,n{+}1\}$. Define the defect
$\delta(n)=a_{\mathrm{B}}(f(n))-n\in\{0,1\}$
and the defect set $\mathcal{D}=\{n\geq 1:\delta(n)=1\}$. The first elements of $\mathcal{D}$ (A395251) are
[…]
Write $\theta_{n}=\{cn+d\}\in[0,1)$, so that $a_{\mathrm{B}}(n)=cn+d-\theta_{n}$.
6.1 Defect positions as return times
We now show that the defect positions are exactly the return times of the Sturmian rotation to an explicit interval, and that every defect position is a repeat of $a_{\mathrm{B}}$.
Theorem 24.
For $d=2(\sqrt{2}{-}1)$ and $n\geq 2$:
$\delta(n)=1\iff\theta_{n-1}\in[1-c,\,1/2].$
Consequently:
  1. (a)
    Every $n\in\mathcal{D}$ is a repeat of $a_{\mathrm{B}}$, i.e. $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$.
  2. (b)
    The set $\mathcal{D}$ has natural density
    $\rho(\mathcal{D})=c-\tfrac{1}{2}=\frac{\sqrt{2}-1}{2}\approx 0.2071.$
Proof.
Using $a_{\mathrm{B}}(n)=cn+d-\theta_{n}$ and $a_{\mathrm{B}}(n{-}1)=c(n{-}1)+d-\theta_{n-1}$:
$f(n)=a_{\mathrm{B}}(n)+a_{\mathrm{B}}(n{-}1)=c(2n{-}1)+2d-\theta_{n}-\theta_{n-1}.$
Since $c^{2}=1/2$:
$cf(n)+d=n-\tfrac{1}{2}+d(2c{+}1)-c(\theta_{n}+\theta_{n-1}).$
At $d=2(\sqrt{2}{-}1)=4c-2$,
$d(2c{+}1)=(4c-2)(2c{+}1)=8c^{2}+2c-2=2,$
so
$cf(n)+d=n+\tfrac{3}{2}-c(\theta_{n}+\theta_{n-1}).$ (33)
Applying the floor, $a_{\mathrm{B}}(f(n))=n+1$ precisely when $n{+}1\leq cf(n)+d<n+2$, which using (33) rearranges to
$\tfrac{1}{2}\leq-c(\theta_{n}+\theta_{n-1})<\tfrac{1}{2},$
equivalently $\theta_{n}+\theta_{n-1}\leq 1/(2c)=c$ (the lower bound is automatic).
Now $\theta_{n}=\{\theta_{n-1}+c\}$. Since $c<1$, two cases arise.
If $\theta_{n-1}<1-c$, then $\theta_{n}=\theta_{n-1}+c$, and
$\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c\leq c$
forces $\theta_{n-1}=0$, which never happens since $d$ is irrational.
If $\theta_{n-1}\geq 1-c$, then $\theta_{n}=\theta_{n-1}+c-1$, and
$\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c-1\leq c$
becomes $\theta_{n-1}\leq 1/2$, yielding $\theta_{n-1}\in[1-c,1/2]$.
This interval is non-empty since 1/2>1c1/2>1-c, with length $1/2-(1-c)=c-1/2=(\sqrt{2}{-}1)/2$. The characterisation is established.
For (a), in the effective case $\theta_{n}=\theta_{n-1}+c-1\in[0,c-1/2]$. Since c1/2<1cc-1/2<1-c, we have $\theta_{n}<1-c$, so $\theta_{n}+c<1$ and $a_{\mathrm{B}}(n{+}1)-a_{\mathrm{B}}(n)=\lfloor\theta_{n}+c\rfloor=0$.
For (b), the sequence $(\theta_{n})_{n\geq 0}$ is the orbit of the irrational rotation $\theta\mapsto\{\theta+c\}$, uniformly equidistributed in $[0,1)$ by Weyl’s theorem. The set of $n$ with $\theta_{n-1}\in[1-c,1/2]$ therefore has natural density equal to the length of that interval. ∎
The absorption mechanism now reads explicitly. At a defect position, the Sturmian address $\theta_{n}$ lies in the low interval [0,c1/2][0,c-1/2], forcing $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$. The identity (5) then survives because the extra +1+1 in $a_{\mathrm{B}}(f(n))$ is absorbed by the outer application of $a_{\mathrm{B}}$,
$a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n).$
6.2 The gap sequence
The characterisation $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in[1-c,1/2]\}$ expresses $\mathcal{D}$ as the return-time set of the rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2]. By the three-distance theorem of Sós [11] (see Alessandri and Berthé [10] for a modern exposition within combinatorics on words), the gaps take at most three values. The return-time analysis of Proposition 25 below identifies these values as $3$, 44, and $7$, and computes the corresponding letter frequencies.
The letter $4$ is always preceded and followed by $3$. Every occurrence of $4$ in the gap sequence appears within a block $3,4,3$, verified numerically for $k\leq 10^{5}$.
Proposition 25.
The gap sequence of $\mathcal{D}$ on the alphabet $\{3,4,7\}$ has letter frequencies
$f_{3}=\sqrt{2}-1,\quad f_{4}=3-2\sqrt{2},\quad f_{7}=\sqrt{2}-1,$
matching Theorem 24(b), and the average gap is $2(\sqrt{2}+1)$. The gap sequence of $\mathcal{D}$ is A395253.
Proof.
The defect set is $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in J\}$ with J=[1c, 1/2]J=[1-c,\,1/2], so $|J|=c-1/2=(\sqrt{2}-1)/2$. For each return time $i\geq 1$, let
$J_{i}=\{\theta\in J:R_{c}^{i}\theta\in J\text{ and }R_{c}^{k}\theta\notin J\text{ for }1\leq k<i\},$
where $R_{c}\colon\theta\mapsto\{\theta+c\}$. We compute the sub-intervals directly from the iterates. Using $c\in(1/2,1)$ and 2c<3c1<3c<22c<3c-1<3c<2, the values $kc\bmod 1$ for small $k$ are
$R_{c}\theta=\theta+c-1,\qquad R_{c}^{2}\theta=\theta+2c-1,\qquad R_{c}^{3}\theta=\theta+3c-2$
for $\theta\in J$. Neither $R_{c}\theta\in[0,c-1/2]$ nor $R_{c}^{2}\theta\in[c,2c-1/2]$ meets $J$, so every return time is at least $3$.
Return time $3$ occurs when $\theta+3c-2\in[1-c,1/2]$, equivalently $\theta\in[3-4c,\,5/2-3c]$. Since 1c>34c1-c>3-4c and 5/23c<1/25/2-3c<1/2, intersection with $J$ gives
$J_{3}=[1-c,\,5/2-3c],\qquad|J_{3}|=(5/2-3c)-(1-c)=3/2-2c=\tfrac{3-2\sqrt{2}}{2}.$
For $\theta\in J\setminus J_{3}=(5/2-3c,\,1/2]$, the return time exceeds $3$. The next iterate $R_{c}^{4}\theta=\theta+4c-3$ lies in $J$ when $\theta\in[4-5c,\,7/2-4c]$. Intersection with (5/23c, 1/2](5/2-3c,\,1/2] gives
$J_{4}=(4-5c,\,1/2],\qquad|J_{4}|=1/2-(4-5c)=5c-7/2=\tfrac{5\sqrt{2}-7}{2}.$
The remainder $J_{7}=(5/2-3c,\,4-5c]=J\setminus(J_{3}\sqcup J_{4})$ has length
$|J_{7}|=(4-5c)-(5/2-3c)=3/2-2c=\tfrac{3-2\sqrt{2}}{2},$
and $R_{c}^{7}\theta\in J$ for every $\theta\in J_{7}$ by direct calculation. The three-distance theorem [11, 10] rules out return times other than $3$, 44, 77, so $J=J_{3}\sqcup J_{4}\sqcup J_{7}$ up to the two measure-zero endpoints.
The orbit $(\theta_{n-1})_{n\geq 2}$ equidistributes in $[0,1)$ by Weyl’s theorem, so each $J_{i}$ is visited with density $|J_{i}|$, and the frequency of gap value $i$ among defect positions is
$f_{i}=\frac{|J_{i}|}{|J|}.$
Substituting the lengths and using $(\sqrt{2}-1)(\sqrt{2}+1)=1$ to rationalise yields
$f_{3}=f_{7}=\sqrt{2}-1,\qquad f_{4}=3-2\sqrt{2}.$
The average gap is
$3f_{3}+4f_{4}+7f_{7}=3(\sqrt{2}{-}1)+4(3{-}2\sqrt{2})+7(\sqrt{2}{-}1)=2\sqrt{2}+2=2(\sqrt{2}{+}1),$
which equals the reciprocal of the density $\rho(\mathcal{D})=(\sqrt{2}-1)/2$ established in Theorem 24(b). ∎
Conjecture 26.
The gap sequence of $\mathcal{D}$ belongs to the minimal subshift generated by the primitive substitution
$\sigma\colon\quad 3\mapsto 3,4,\qquad 4\mapsto 3,7,\qquad 7\mapsto 3,7,7.$ (34)
The substitution has incidence matrix
$M=\begin{pmatrix}1&1&1\\ 1&0&0\\ 0&1&2\end{pmatrix}$,
characteristic polynomial $\lambda^{3}-3\lambda^{2}+\lambda+1=(\lambda-1)(\lambda^{2}-2\lambda-1)$, and Perron eigenvalue $1+\sqrt{2}$ (the silver ratio). Since $M^{2}$ has all positive entries, $\sigma$ is primitive of index $2$, and the Perron eigenvector gives letter frequencies matching Proposition 25.
The conjecture is supported by numerical evidence (verified for $k\leq 10^{5}$) and by the following natural candidate proof, which we leave for future work to complete.
Sketch of a Rauzy induction argument. Using the sub-intervals $J_{3}$, $J_{4}$, $J_{7}$ computed in the proof of Proposition 25, the first-return map $R_{c}^{\mathrm{ret}}\colon J\to J$ permutes the three pieces. Direct inspection of the images $R_{c}^{3}(J_{3})$, $R_{c}^{4}(J_{4})$, $R_{c}^{7}(J_{7})$ inside $J$ suggests the transitions
$J_{3}\to J_{3}J_{4},\qquad J_{4}\to J_{3}J_{7},\qquad J_{7}\to J_{3}J_{7}J_{7},$
which transpose to (34). A complete proof would require the full Rauzy induction machinery on three letters, see [7, Chapter 6].
Remark 27.
The three gap values $3,4,7$ are exactly $q_{2}$, $q_{3}-q_{2}$, and $q_{3}$ in the Pell sequence, and the conjectural Perron eigenvalue $1+\sqrt{2}$ of $\sigma$ is the dominant root of the Pell recurrence $X_{n+1}=2X_{n}+X_{n-1}$. Both facts instantiate the Pell–Ostrowski structure described in Section 4.5.
6.3 A two-letter coarsening
Grouping the three gap values into two classes simplifies the structure. Encoding $3,4$ as a short letter $S$ and $7$ as a long letter $L$ yields a two-letter sequence with frequencies $2-\sqrt{2}$ for $S$ and $\sqrt{2}-1$ for $L$. Both frequencies lie in $\mathbb{Q}(\sqrt{2})$, consistent with $a_{\mathrm{B}}$ being a rotation sequence of slope $c=1/\!\sqrt{2}$. Empirically the $S/L$ sequence appears to be Sturmian of slope $\sqrt{2}-1$, but a rigorous identification is left as an open problem.
7 Open problems
  1. 1.
    Is the affine solution (7) the unique continuous monotone solution of the continuous equation (6)? By Theorem 4, any such solution is a homeomorphism satisfying $\varphi^{-1}(x)=\sum_{j=0}^{r-1}\varphi(x{-}j)$.
  2. 2.
    Among nondecreasing solutions of (1) for $r\geq 3$, classify all monotone solutions. For $r=2$ the tree structure is described in [3]. The general case is open.
  3. 3.
    In the Ostrowski numeration system for $c=[0;1,2,2,2,\ldots]$ (see [9]), with convergents $p_{k}/q_{k}$ and digits $0\leq d_{k}\leq a_{k+1}$, verify that $a_{\mathrm{B}}(n)=\sum d_{k}p_{k}$ whenever $n=\sum d_{k}q_{k}$. This digit-swap property has been confirmed numerically for $n\leq 10^{6}$ and would give a carry-free interpretation of the identity $a_{\mathrm{B}}(a_{\mathrm{B}}(n)+a_{\mathrm{B}}(n{-}1))=n$.
  4. 4.
    Identify the analogue of Conjecture 26 for each $r\geq 3$. For $r=3$ the defect set at the right endpoint $d=(5\sqrt{3}{-}3)/6$ begins
    $16,\,23,\,42,\,49,\,61,\,68,\,87,\,94,\ldots$
    (A395252), with gap values $\{7,\,12,\,19\}$ corresponding to the consecutive convergent denominators $7$ and $19$ of $1/\!\sqrt{3}$ and their difference. A primitive substitution on $\{7,12,19\}$ should govern the structure, with Perron eigenvalue related to $\sqrt{3}$.